### v1

# Facial Expression Analysis and Its Potentials in IoT Systems: A Contemporary Survey

Abstract.
Facial expressions convey human emotions and can be categorized into macro-expressions (MaEs) and micro-expressions (MiEs) based on duration and intensity. While MaEs are voluntary and easily recognized, MiEs are involuntary, rapid, and can reveal concealed emotions. The integration of facial expression analysis with Internet-of-Thing (IoT) systems has significant potential across diverse scenarios. IoT-enhanced MaE analysis enables real-time monitoring of patient emotions, facilitating improved mental health care in smart healthcare. Similarly, IoT-based MiE detection enhances surveillance accuracy and threat detection in smart security. This work aims at providing a comprehensive overview of research progress in facial expression analysis and explores its integration with IoT systems. We discuss the distinctions between our work and existing surveys, elaborate on advancements in MaE and MiE techniques across various learning paradigms, and examine their potential applications in IoT. We highlight challenges and future directions for the convergence of facial expression-based technologies and IoT systems, aiming to foster innovation in this domain. By presenting recent developments and practical applications, this study offers a systematic understanding of how facial expression analysis can enhance IoT systems in healthcare, security, and beyond.
Facial expression analysis, Internet of Thing, macro- and micro-expressions
Nomenclature
Abbreviations Definitions
AFEW 7.0 The acted facial expressions in the wild
AU Action unit
BERT Bidirectional encoder representations from transformers
BU-3DFE Binghamton university 3D facial expression
CAER Context-aware emotion recognition
CASME Chinese Academy of Sciences micro-expression
CASME II Chinese Academy of Sciences micro-expression II
CAS(ME)2 Chinese Academy of Sciences macro-expressions and micro-expressions
CK+ The Extended Cohn-Kanade Dataset
CNN Convolutional neural network
DFEW Dynamic facial expression in the wild
ExpW Expression in-the-wild
FACS Facial action coding system
FER-2013 Facial expression recognition 2013
GAN Generative adversarial network
HOG Histogram of gradient
ISED Indian spontaneous expression database
IoT Internet of Things
JAFFE Japanese female facial expression
KDEF Karolinska directed emotional faces
LBP Local binary pattern
LBP-TOP Local binary pattern from three orthogonal planes
LSTM Long short-term memory
MaE Macro-expression
MDMO Main directional mean optical-flow
MEVIEW Micro-expression videos in the wild
MiE Micro-expression
MMEW Micro-and-macro in expression warehouse
PEDFE Padova emotional dataset of facial expressions
RaFD Radboud faces database
RNN Recurrent neural network
ROI Region of interest
SAMM Spontaneous actions and micro-movements
SFEW 2.0 Static facial expressions in the wild
SMIC Spontaneous micro-expression corpus
TFD The Toronto face database
UAR Mean class accuracy
ViT Vision transformer
WAR Overall accuracy
1. Introduction
Since its inception in 1999, Internet-of-Things (IoT) technology has seamlessly integrated into the fabric of modern society, driving key innovations in both civilian and industrial sectors. The widespread adoption of IoT technology generates skyrocketing amounts of daily data that were traditionally processed in clusters of cloud servers. Due to the ever-increasing privacy concerns, the information process is shifted from the cloud to the network edge. The low access latency and high data security of edge computing enable new application domains, such as emotion detection, traffic surveillance, and healthcare (Dong et al., 2019, 2018).
Figure 1. A framework of edge-driven facial expression analysis.
Figure 2. An illustration of MaEs and MiEs  in micro-and-macro in expression warehouse (Ben et al., 2021).
Since facial expressions convey 55% of emotional messages during the exchange of moods, thoughts, feelings, and mental states, they have become a medium for sensing emotions in front of facility screens (Mehrabian and Wiener, 1967). Recent advancements in IoT technologies have driven the widespread application of facial expression analysis in practical scenarios. As shown in Fig. 1, IoT devices can first collect raw facial expression data for pre-processing. To fully leverage low access latency and high data security during the implementation of real-time privacy-preserving systems, the pre-processed features are then uploaded to edge devices for decision-making. Facial expressions can be divided into macro-expressions (MaEs) and micro-expressions (MiEs) based on their duration and intensity. Two types of emotions within MaEs and MiEs are illustrated in Fig.2. In terms of duration, MaEs approximately last 0.5 to 4 seconds, while MiEs last less than 0.5 seconds (Shen et al., 2012). The short duration of MiEs conveys subtle emotional message, which makes emotion perception from MiEs a challenging task without specialized techniques and training. The major differences between MaEs and MiEs are described as follows.
  • MaEs can be perceived during regular interactions. The recognition accuracy of MaEs  can exceed 97% in a controlled laboratory due to the spontaneity and noise immunity. The MaEs can be applied in multiple scenarios, e.g., autonomous driving, psychological health assessment, and human-computer interaction.
  • MiEs are involuntary and rapid. Due to the short duration and subtle emotional cues, the detection of MiEs  can be quite challenging with the naked eye. It is reported that the average recognition accuracy of MiEs is around 47% even after specialized training (Frank et al., 2009). Nevertheless, MiEs can reveal the true emotions of a person since they are instinctive and cannot be concealed. Hence, numerous applications of MiEs in IoT systems are being explored, such as lie detection, criminal identification, and security control.
The objective of facial expression analysis lies in detecting the emotions of humans from MaEs and MiEs. The research works on MaE analysis are mainly based on deep models that detect human emotions by leveraging the isolated static images and sequential dynamic images of MaEs. More specifically, the deep static MaE analysis performs well for two scenarios: (1) when several isolated images are available; and (2) when real-time expression recognition is required. The deep dynamic MaE analysis can benefit from the comprehensive understanding of temporal dynamics of MaEs. Since the emotion recognition of MiEs is challenging due to the short duration and subtle emotional cues, the research works on holographic MiE analysis consist of two parts, i.e., MiE spotting and MiE recognition. The MiE spotting aims at detecting MiEs within a given video, and the MiE recognition then uses the detected MiEs for MiE classification within a set of MiE categories. Different from the MaE analysis, the MiE analysis can provide valuable insights into potentially hidden human emotions. In this work, we aim at providing a contemporary tutorial on facial expression analysis that ranges from the used datasets to the detailed technical methods. Compared with the previous surveys on facial expression analysis, our review covers a broader perspective on both MaE and MiE analysis. The emerging techniques on MaE and MiE analysis are discussed on deep MaE recognition and holographic MiE analysis. Besides, we also discuss the potential integrations of MaE and MiE analysis with the IoT systems in order to inform new directions of facial expression application.
As shown in Fig. 3, the remaining work is organized as follows. Section 2 introduces the recent surveys on MaE and MiE analysis and major differences from our work. Section 3 presents the datasets for MaE and MiE analysis. Section 4 discusses the deep MaE recognition. Section 5 discusses the holographic MiE recognition that includes spotting and recognition procedures. Section 6 discusses the applications of MaE and MiE analysis in IoT systems. Section 7 discusses the challenges and future directions. The conclusions are presented in Section 8.
Figure 3. The organization of the remaining work.
2. Recent Surveys on Facial Expression Analysis
Facial expression analysis is a critical tool for understanding human emotions and intentions. The research on facial expression analysis started in the 1970s (Ekman and Friesen, 1978, 1971). Over the years of development, facial expression analysis has penetrated into our human daily lives. More specifically, facial expression analysis can be categorized into MaE analysis and MiE analysis. Therefore, we are motivated to review recent research contributions on both MaE and MiE analysis to better understand the current research stage.
2.1. Recent Surveys on MaE Analysis
MaE analysis is a burgeoning field at the intersection of computer vision and human emotion recognition and can offer transformative potential in healthcare, security, and human-computer interaction. Pioneering research has laid the groundwork for recognizing subtle expressions. For example, Pantic et al. (Rothkrantz and Pantic, 2000) introduced three key problems in automatic MaE analysis: (1) detection of an image segment as a face; (2) extraction of facial expression information; and (3) classification of the expression. Yet, significant challenges remain in achieving robust, scalable, and generalizable solutions to MaE Analysis. Leveraging the power of deep learning, researchers endeavor to address these challenges, advancing the field through innovative algorithms and practical IoT applications. In (Li and Deng, 2020), Li et al. provided a comprehensive review of MaE recognition based on deep learning. More specifically, the available datasets for MaE analysis and the principles of data selection and evaluation were explored (Li and Deng, 2020). Later, Canal et al. (Canal et al., 2022) conducted an exhaustive systematic literature review encompassing $94$ leading methods from $51$ related articles. Their review meticulously delineates the computational workflow for MaE recognition, such as, preprocessing, feature extraction, and classification algorithms. Through a rigorous statistical analysis, Canal et al. (Canal et al., 2022) revealed that traditional machine learning methods suffer from limited generalization capabilities though the high precision in classification can be achieved. With the powerful feature capabilities, deep learning techniques thrive in scenarios involving larger and more diverse datasets. Karnati et al. (Karnati et al., 2023) introduced the cutting-edge models that are tailored for deep MaE recognition across diverse input modalities. The comprehensive evaluations provide compelling evidence of enhanced performance and generalization potential of deep models that set new benchmarks for future research (Karnati et al., 2023).
Over the last decades, research on MaE analysis has undergone remarkable evolution that was transitioned from a formidable challenge to a structured and highly efficient process. Once fraught with complexities, the MaE analysis workflow now adheres to a robust standard pipeline that encompasses preprocessing, feature extraction, and classification. The integration of deep learning into this pipeline has been a transformative milestone, propelling algorithmic performance to exceed 97% accuracy under controlled laboratory conditions—an unprecedented achievement. Despite these advancements, comprehensive overviews remain scarce to bridge the gap between foundational research and practical applications. This limitation hampers the widespread adoption of MaE technologies in real-world (a.k.a., wild) scenarios, especially in areas like IoT integration and multimodal recognition. By bridging the gap between foundational research and wild application, we first introduce deep MaE recognition algorithms in Section 4 and the MaE applications based on IoT in Section 6.1. Our analysis not only consolidates existing knowledge but also sets the stage for pioneering applications, ensuring our research addresses the pressing needs of this rapidly advancing field.
2.2. Recent Surveys on MiE Analysis
As MiE analysis continues to gain significant attention across academia and industry, numerous surveys have systematically reviewed advancements in MiE analysis from diverse perspectives. For example, Ben et al. (Ben et al., 2021) conducted the first comprehensive survey that offers a systematic examination of MiE analysis within a unified evaluation framework. A neuropsychological perspective was provided on the distinction between MiEs and MaEs, i.e., the greater inhibitory effect of MiE on facial expressions in (Ben et al., 2021). Additionally, several representative MiE datasets were also introduced and could be categorized into four distinct types: (1) instructing participants to perform specific MiEs (Shreve et al., 2011; Polikovsky et al., 2009); (2) constructing high-stakes scenarios (Warren et al., 2009); (3) eliciting emotions through video stimuli while maintaining neutral expressions (Li et al., 2013; Yan et al., 2013, 2014; Davison et al., 2016; Qu et al., 2017); and (4) capturing real high-stakes situations (Husák et al., 2017). Notably, the micro-and-macro in expression warehouse (MMEW) dataset that contains high-resolution samples with a balanced distribution of MaE and MiE was proposed in (Ben et al., 2021). Xie et al. (Xie et al., 2022) further advanced the understanding of MiE recognition by delving into critical areas, such as, macro-to-micro adaptation, recognition based on apex frames, and analysis leveraging facial action units (AUs). Their detailed exploration offers a more nuanced perspective, complementing prior works and addressing emerging challenges in the field. Li et al. (Li et al., 2022d) presented a pioneering survey on MiE analysis through the lens of deep learning techniques. They systematically reviewed existing deep learning approaches by analyzing datasets, delineating the stepwise pipelines for MiE recognition, and conducting performance comparisons with state-of-the-art methods. Their survey introduced a novel taxonomy for deep learning-based MiE recognition that can classify input data into static, dynamic, and hybrid categories. Furthermore, Li et al. (Li et al., 2022d) meticulously examined the architectural components of deep learning models, including network blocks, structural designs, training strategies, and loss functions. Their findings emphasized that leveraging multiple spatiotemporal features as input yields superior performance in MiE recognition. Zhao et al. (Zhao et al., 2023a) provided a holistic overview of MiE research from foundational psychological studies and early computer vision efforts to advanced computational MiE analysis. Zhao et al. (Zhao et al., 2023a) highlighted key research directions that include MiE AU detection and MiE generation, and explored practical applications (e.g., covert emotion recognition and professional training). Additionally, Zhao et al. (Zhao et al., 2023a) identified pressing challenges in wild MiE applications, such as, data privacy, data protection, fairness, diversity, and the regulated use of MiE technologies.
Previous surveys on MiE analysis have comprehensively examined the field from various perspectives, e.g., MiE analysis, dataset creation, and psychological foundations of MiE development. However, critical aspects such as advanced MiE analysis and practical MiE applications remain underexplored. To address the aforementioned gaps, our work delves into state-of-the-art MiE analysis in Section 5. Additionally, we explore the transformative potential of MiE applications in the IoT system in Section 6.2. We expect to provide a holistic perspective that can bridge foundational research with cutting-edge advancements and wild applicability.
3. Datasets for Facial Expression Analysis
Facial expression analysis relies on high-quality datasets to drive advancements in model development and evaluation. In this section, we will provide a holistic introduction to the current datasets used for MaE and MiE analysis.
3.1. MaE Datasets
The recent MaE datasets including Japanese female facial expression (JAFFE) (Lyons et al., 1998), Karolinska directed emotional faces (KDEF) (Goeleven et al., 2008), MMI (Valstar et al., 2010; Pantic et al., 2005), Binghamton university 3D facial expression (BU-3DFE) (Yin et al., 2006), the Toronto face database (TFD) (Susskind et al., 2010), The Extended Cohn-Kanade Dataset (CK+) (Lucey et al., 2010), Radboud faces database (RaFD) (Langner et al., 2010), MUG (Aifanti et al., 2010), Multi-PIE (Gross et al., 2010), static facial expressions in the wild (SFEW 2.0) (Dhall et al., 2011), Oulu-CASIA (Zhao et al., 2011), facial expression recognition 2013 (FER-2013) (Goodfellow et al., 2013), Indian spontaneous expression database (ISED) (Happy et al., 2017), EmotioNet (Fabian Benitez-Quiroz et al., 2016), wild affective faces database (RAF-DB) (Li et al., 2017a), AffectNet (Mollahosseini et al., 2019), the acted facial expressions in the wild (AFEW 7.0) (Dhall et al., 2017), expression in-the-wild (ExpW) (Zhang et al., 2018b), Aff-Wild2 (Kollias and Zafeiriou, 2019), context-aware emotion recognition (CAER) (Lee et al., 2019), dynamic facial expression in the wild (DFEW) (Jiang et al., 2020), FERV39k (Wang et al., 2022c), Padova emotional dataset of facial expressions (PEDFE) (Miolla et al., 2023), CalD3r + MenD3s (Ulrich et al., 2024). Table 1 illustrates the detailed information of the MaE datasets.
Table 1. An Overview of MaE Datasets
Dataset Subjects MaE samples Annotation Collection Details
Num Ethnicities Num
Resolution/
Frame rate
Emotion classes FACS AU Index Condittion Elicitation Year
JAFFE (Lyons et al., 1998) 10 1 213 images 256 × 256 BEs N/A N/A N/A Lab P 1998
KDEF (Goeleven et al., 2008) 70 N/A 4900 images 562 × 762 BEs N/A N/A NA Lab P & S 1998
MMI (Valstar et al., 2010; Pantic et al., 2005) 25 3
704 images &
2900 sequences
720 × 576 BEs Yes Yes NAF Lab P 2005
BU-3DFE (Yin et al., 2006) 100 6
2500 3D
images
1040 × 1329 BEs N/A N/A N/A Lab P 2006
TFD (Susskind et al., 2010) N/A N/A 112,234 images 48 × 48 BEs N/A N/A N/A Lab P 2010
CK+ (Lucey et al., 2010) 123 3 593 sequences 640 × 490 BEs Yes Yes NA Lab P & S 2010
RaFD (Langner et al., 2010) 67 1 1608 images 256 × 256
BEs plus contempt
without neutral
N/A Yes N/A Lab P 2010
MUG (Aifanti et al., 2010) 52 1 1462 sequences 896 × 896, 16fps BEs N/A N/A NAF Lab P 2010
Multi-PIE (Gross et al., 2010) 337 4 755,370 images 3072 × 2048
smile, surprise, squint,
disgust, scream and neutral
N/A N/A N/A Lsb P 2010
SFEW 2.0 (Dhall et al., 2011) N/A N/A 1766 images Multiple BES N/A N/A N/A Movie P & S 2011
Oulu-CASIA (Zhao et al., 2011) 36 2
1920 images &
2880 sequences
320 × 240 BEs without neutral N/A N/A NA Lab P 2011
FER-2013 (Goodfellow et al., 2013) N/A N/A 35762 imgaes 48 × 48 BEs N/A N/A N/A Internet P & S 2012
ISED (Happy et al., 2017) 50 1 428 sequences 1920 × 1080, 50fps
sadness, surprise, happiness
and disgust
N/A N/A N/A Lab S 2015
EmotioNet (Fabian Benitez-Quiroz et al., 2016) N/A N/A 1,000,000 images Multiple
23 basic expressions or
compound expressions
N/A Yes N/A Internet P & S 2016
RAF-DB (Li et al., 2017a) N/A N/A 29,672 images 768 × 768
BEs and twelve
compound expressions
N/A N/A N/A Internet P & S 2016
AffectNet (Mollahosseini et al., 2019) N/A N/A 450,000 images Multiple
BEs, arousal
and valence
N/A N/A N/A Internet P & S 2017
AFEW 7.0 (Dhall et al., 2017) N/A N/A 1809 sequences Multiple BEs N/A N/A N/A Movie P & S 2017
ExpW (Zhang et al., 2018b) N/A N/A 91,793 images Multiple BEs N/A N/A N/A Internet P & S 2018
Aff-Wild2 (Kollias and Zafeiriou, 2019) 458 N/A 558 sequences Multiple
BEs, arousal
and valence
N/A Yes N/A Internet P & S 2019
CAER (Lee et al., 2019) N/A N/A 13,201 sequences Multiple BEs N/A N/A N/A TV show P & S 2019
DFEW (Jiang et al., 2020) N/A N/A 16,372 sequences Multiple BEs N/A N/A N/A Movie P & S 2020
FERV39k (Wang et al., 2022c) N/A N/A 38,935 sequences 3840 × 2160 BEs N/A N/A N/A Internet P & S 2021
PEDFE (Miolla et al., 2023) 56 N/A 1458 sequences 1920 × 1080, 30fps BEs without neutral N/A N/A NA Lab P & S 2022
CalD3r + MenD3s (Ulrich et al., 2024) 104 + 92 1+1
4,678 images
4,038 images
640 × 480, 30fps BEs N/A N/A N/A Lab S 2024
  • 1 BEs: Basic expressions (seven); N/A: Not applicable; P: Posed; S: Spontaneous;
  • 2 NAF: Onset frame, Apex frame, Offset frame, respectively.
The basic expressions of most MaE datasets can be divided into seven categories, i.e., anger, neutral, disgust, fear, happiness, sadness, and surprise. Although being considered as a basic emotion of the face, contempt expression is not common in most MaE datasets. Few datasets (e.g., CK+ and RaFD) have collected the contempt expression; therefore, the number of samples on contempt expression is limited. In the CK+ dataset, the number of contempt expression samples accounts for 5% of total samples. In addition to the seven basic expressions in MaE datasets, several datasets (e.g., AffectNet and AFF-wild2) also used the continuous-valued valence and arousal to describe the intensity of expression. Moreover, only the ExpW dataset considered the compound expressions. In the ExpW dataset, 23 basic or compound emotion categories were used for emotion classification.
The early datasets were usually collected from the laboratory and required a certain number of subjects to join. The collection of these datasets required a lot of labor, which may include subjects participating in the collection, guides of collection experiments, and professional annotators. In addition to the annotation of expression categories, several datasets (e.g., CK+, MMI, and Oulu-CASIA) required additional information in terms of facial action coding system (FACS), AU, and index. Typically, CK+ and Oulu-CASIA have utilized the annotation information of the index to label their sequences with the onset and peak of expressions. MMI and CK+ have labeled the FACS and AU to provide more additional information. In addition to the labor-consuming collection of MaE datasets, EmotioNet has utilized the automatic annotation algorithm to label the data collected from the internet. In this dataset, a total of 9500,000 images were annotated with AUs, AU intensities, and emotion categories.
The early MaE datasets (e.g., JAFFE) were utilized to obtain the posed images and sequences. These datasets contained the posed data for the front view. To meet the requirement of practical MaE in wild conditions, many datasets (e.g., Multi-PIE, RaFD) provided the MaE data with various environmental conditions including multiple head poses, occlusions, and illuminations. Typically, the Multi-PIE contained MaE images under 19 illumination and 15 viewpoint conditions in four sessions. In their process of dataset collection, each subject was recorded between $-90^{\circ}$ to $90^{\circ}$ with an interval of $15^{\circ}$. MaE in RaFD was recorded at the same moment through five different camera angles and shown with three different gaze directions. In addition, the MaE datasets collected from the internet (e.g., AffectNet and ExpW) contained an amount of MaE data with more wild scenarios. These datasets can benefit the robustness and generalization of further MaE research.
Table 2. Spontaneous MiE Datasets
Dataset Subjects MiE Videos Annotations
Num Mean age Ethnicities Num Resolution Frame rate Emotion classes FACS AU Index
SMIC (Li et al., 2013) HS 16 N/A 3 164 640$\times$480 100 Pos (51) Neg (70) Sur (43) N/A N/A NF
VIS 8 71 25 Pos (28) Neg (23) Sur (20) N/A
NIR 8 71 25 Pos (28) Neg (24) Sur (19) N/A
CASME (Yan et al., 2013) 35 22.03 1 195
640$\times$480
1280$\times$720
60
Hap (5) Dis (88) Sad (6) Con (3)
Fea (2) Anx (28) Sur (20) Rep (40)
Yes Yes NAF
CASME II (Yan et al., 2014) 35 22.03 1 247 640$\times$480 200
Hap (33) Sur (25) Dis (60)
Rep (27) Oth (102)
Yes Yes NAF
CAS(ME)2 (Qu et al., 2017) 22 22.59 1 57 640$\times$480 30 Hap (51) Neg (70) Sur (43) Oth (19) Yes Yes NAF
SAMM (Davison et al., 2016) 32 33.24 13 159 2040$\times$1088 200
Hap (24) Dis (8) Sad (3) Fea (7)
Sur (13) Ang (20) Oth (84)
Yes Yes NAF
MEVIEW (Husák et al., 2017) 16 N/A N/A 40 1280$\times$720 25
Hap (6) Dis (1) Fea (3) Sur (9)
Con (6) Ang (2) Conf (13)
Yes Yes NF
CAS(ME)3 (Li et al., 2022b) 247 22.74 1 1059 1280$\times$720 30
Hap (992) Dis (2528) Sad (635) Fea (892)
Sur (1208) Con (401) Ang (20) Oth (251)
Yes Yes NAF
MMEW (Ben et al., 2021) 36 N/A 1 16 1920$\times$1080 25
Hap (36) Dis (72) Sad (13) Fea (16)
Sur (80) Ang (8) Oth (84)
Yes Yes NAF
4DME (Li et al., 2022a) 56 27.8 3 1068
640$\times$480
1200$\times$1600
30/60
60
Pos (34) Neg (127) Sur(30) Rep(6)
PS(13) NS(8) RS (3) PR(8) NR (7) Oth(31)
Yes Yes NAF
  • 1Pos: Positivity; Neg: negativity; Sur: Surprise; Hap: Happiness; Dis: Disgust; Sad: Sadness; Con: Contempt; Fea: Fear; Anx: Anxiety; Rep: Repression; Ang: Anger; Conf: Confusion; Oth: Others; PS: Pos w/ Sur; NS: Neg w/ Sur; RS: Rep w/ Sur; PR: Pos w/ Rep; NR: Neg w/ Rep;
  • 2N/A: Not applicable; NAF: Onset frame, Apex frame, and Offset frame, respectively.
3.2. MiE Datasets
In recent years, multiple datasets have been built for further development of MiE analysis. Initially, researchers collected the MiE databases (e.g., Polikvsky’s database (Polikovsky et al., 2009) and USF-HD (Shreve et al., 2011)) by asking the participants to pose or mimic facial expressions. These posed datasets supported the preliminary studies of MiE, but were not used for current analysis due to the properties of private datasets and not capturing the nature of MiE. Later, more spontaneous datasets, including the spontaneous micro-expression corpus (SMIC) (Li et al., 2013), Chinese Academy of Sciences micro-expression (CASME) (Yan et al., 2013), Chinese Academy of Sciences micro-expression II (CASME II) (Yan et al., 2014), Chinese Academy of Sciences macro-expressions and micro-expressions (CAS(ME)2) (Qu et al., 2017), spontaneous actions and micro-movements (SAMM) (Davison et al., 2016), micro-expression videos in the wild (MEVIEW) (Husák et al., 2017), CAS(ME)3 (Li et al., 2022b), MMEW (Ben et al., 2021), and 4D spontaneous micro expression database (4DME) (Li et al., 2022a), were utilized in MiE research. The detailed information on the spontaneous MiE datasets is provided in Table 2.
Due to the subtle and fleeting nature of MiE, the annotation of MiE datasets is a challenging and demanding task that requires a significant amount of time and labor. Coders who label MiE need FACS training and half an hour to detect suitable clips (Pantic, 2009). The labeling of MiE includes the AU and emotion classification. By incorporating the AU into certain emotion taxonomy, the technique of MiE synthesis can be used to generate new samples. The classification of MiE in the dataset is different, which results in inconsistent annotations between various datasets. For example, the classification of MiE in SMIC is positive, negative, and surprising. The classification of MiE in other datasets, such as CASME, MEVIEW, and SAMM, can be divided into more than six categories.
To induce spontaneous MiE, many studies have used a collection paradigm that requires participants to maintain a poker face while watching intensely emotional video clips. MiEs are revealed and recorded using a high-speed camera. This effective and simplified method was applied to five datasets: SMIC, CASME, CASME II, SAMM, and CAS(ME)2. However, this paradigm was criticized for its laboratory setting, which lacks practical or realistic situations. Consequently, studies expanded to more realistic settings. In real-life scenarios, such as high-stakes poker games or TV interviews, participants often conceal their true emotions, leading to MiE occurrences. For example, MEVIEW constructed an in-the-wild MiE dataset from real poker games and TV interviews. Despite its benefits for real-scene analysis, MEVIEW has limited data samples and frequent face pose changes, resulting in side views and occlusions. Subsequently, CAS(ME)3 adopted a trade-off method using mock crime paradigms to collect MiE samples in practical scenes with controllable factors.
Figure 4. Samples of current MiE datasets. These are as follows: (a) SMIC (Li et al., 2013), (b) CASME (Yan et al., 2013), (c) CASME II (Yan et al., 2014), (d) CAS(ME)2 (Qu et al., 2017), (e) SAMM (Davison et al., 2016), (f) MEVIEW (Husák et al., 2017), (g) CAS(ME)3 (Li et al., 2022b), (h) MMEW (Ben et al., 2021), and (i) 4DME (Li et al., 2022a).
[…]
This figure provides the samples of current MiE datasets.
Early spontaneous MiE datasets collected the 2D facial video, which had the advantages of convenient collection and control. With the concern about adequate and accurate MiE analysis, more modalities of the dataset were emphasized. Except for using 2D facial video, the dataset of SMIC implemented the 2D near infrared videos to alleviate the influence of illumination. More recently, the depth information has proven its effectiveness from face recognition (Mian et al., 2007) to expression recognition (Corneanu et al., 2016). A standard RGB image can be constructed into a 3D model of the human face with depth information, which can enhance the robust perception of expression features and cognitive behavior. Therefore, the modal of depth information was applied in the datasets of CAS(ME)3 and 4DME. Besides, CAS(ME)3 enriched the data information with physiological signals and voice signals. Also, 4DME developed more facial multi-modal videos, consisting of reconstructed dynamic 3D facial meshes, grayscale 2D frontal facial videos, Kinect-color videos, and Kinect-depth videos. Fig. 4 has shown the samples in spontaneous MiE datasets.
In addition, some datasets considered the composition of MiE and other expressions for the real-scene situation analysis. For example, MMEW, CAS(ME)2, CAS(ME)3, and 4DME contained the MiE and MaE, which can be employed for detecting MiE from complete videos and further analysis about the evolution of different expressions. Due to the short length of video data, the analysis of MiE was constrained. To collect long videos, several datasets, including CASME, CASME II, and SMIC, were expanded to incorporate frames that do not exhibit MiE before and after the annotated MiE samples. Long MiE videos can benefit more sharper MiE algorithms in more complex situations, such as head movement and verbal behavior that occurred in common scenarios.
Table 3. An Overview of Representative Methods on Deep Static MaE Recognition
Learning
Paradigms
Method Year Pre-processing Block Dataset Performance
Network
Structure
Protocol
Ensemble Learning Kim et al. (Kim et al., 2015) 2016 IN+FA+CE N/A SFEW 2.0 61.6% CNN Hold out
Pons et.al . (Pons and Masip, 2018b) 2018 IN+FA N/A SFEW 2.0 60.5% CNN LOSO
Saurav et al. (Saurav et al., 2022) 2022 IN+FA+CE N/A FER2013/RAF-DB /CK+ /SFEW 2.0 72.77% /86.07% /98.54% /60.5% CNN LOSO
Shao et al. (Shao and Qian, 2019) 2019 FD+FA+DA RES CK+/BU-3DFE/FER2013 86.50%/86.50%/71.14% CNN 10 fold
Reddy et al. (Viswanatha Reddy et al., 2020) 2020 DA N/A AffectNet 59% CNN Hold out
Sharifnejad et al. (Sharifnejad et al., 2021) 2021 FD N/A CK+ 95.33% N/A 10 fold
Hariri et al. (Hariri and Farah, 2021) 2021 N/A N/A BU-3DFE 94.73% CNN 10 fold
Mohan et al. (Mohan et al., 2021) 2021 DA N/A CK+/JAFFE/KDEF/FER2013/RAFDB 98%/98%/96%/78%/83% 2sCNN 10 fold
Xie et al. (Xie and Hu, 2019) 2019 FD N/A CK+/JAFFE 93.46%/94.75% 2sCNN 10 fold
Wadhawan et al. (Wadhawan and Gandhi, 2023) 2023 FD+CE N/A CK+/JAFFE/SFEW 2.0 97.31%/97.14%/44.50% CNN 10 fold+LOSO
Hua et al. (Hua et al., 2019) 2019 DA N/A FER2013/JAFFE/AffectNet 71.91%/96.44%/62.11% 3sCNN 10 fold
Transfer Learning Ng et al. (Ng et al., 2015) 2015 FD N/A Emotion Recognition In The Wild Challenge 55.6% CNN 10 fold
Akhand et al. (Akhand et al., 2021) 2021 FD+DA N/A KDEF/JAFFE 96.51%/99.52% CNN Hold out
Ngo et al. (Ngo and Yoon, 2020) 2020 DA RES AffectNet 60.70% CNN Hold out
Atabansi et al. (Atabansi et al., 2021) 2021 FD+DA N/A Oulu-CASIA 98.11% CNN Hold out
Multi-task Learning Pons et al. (Pons and Masip, 2018a) 2018 DA RES SFEW 2.0 45.9% CNN Hold out
Chen et al. (Chen et al., 2021) 2021 N/A RES AffectNet/RAF 61.98%/87.27% CNN 10 fold
Yu et al. (Yu and Xu, 2022) 2022 DA+FA RES+Atten RAF-DB/SFEW 2.0/CK+/Oulu-CASIA 90.36%/45.78%/98.33%87.32% CNN 10 fold+LOSO
Liu et al. (Liu et al., 2024) 2023 N/A RES+Trans AffectNet/Aff-Wild2 65.80%/64.13 CNN Hold out
Yu et al. (Yu et al., 2020a) 2020 FA+ROI N/A CK+/Oulu-CASIA 99.08%/90.40% GRU 10 fold
Xiao et al. (Xiao et al., 2023) 2023 FD N/A CK+/MMI/RAF-DB 9.07%/84.62%/87.52% CNN 5 fold+Hold out
Zhao et al. (Zhao et al., 2021b) 2021 FA+DA N/A CK+/Oulu-CASIA/MMI 97.85%/89.23%/75.32% CNN 10 fold+Hold out
Chen et al. (Chen et al., 2022b) 2022 FD+FA Atten Multi-PIE/KDEF 88.41%/89.04% CNN 5 fold
Pan et al. (Pan et al., 2022) 2022 DA RES RAF-DB/FER2013 88.30%/68.54% CNN Hold out
Qin et al. (Qin et al., 2024) 2023 DA Trans +Atten RAF-DB 90.97% CNN Hold out
Foggia et al. (Foggia et al., 2023) 2023 DA RES RAF-DB 85.30% CNN 5 fold
Attention-Based Learning Fernandez et al. (Marrero Fernandez et al., 2019) 2019 DA RES+Atten CK+/BU-3DFE 90.30%/82.11% CNN Cross-validation
Li et al. (Li et al., 2020a) 2020 FD RES+Atten CK+/JAFFE/Oulu-CASIA/FER2013 98.68%/98.52%/94.63%/75.82% 2sCNN 5 fold
Li et al. (Li et al., 2019b) 2018 FD+FA RES+Atten RAF-DB/AffectNet 85.07%/58.78% CNN 10 fold
Zhao et al. (Zhao et al., 2021a) 2021 FA+DA RES+Atten CAER/AffectNet/RAF-DB/SFEW 2.0 88.42%,/64.53%,/88.40%,/59.40% CNN Cross-validation
Liu et al. (Liu et al., 2022a) 2022 DA+ROI RES+Atten
RAF-DB/AffectNet/ SFEW 2.0/FER-2013
/AffectNet-8
89.25%/64.54%/61.17%/74.48%
/61.74%
CNN Hold out
Aouayeb et al. (Aouayeb et al., 2021) 2021 DA RES+Trans CK+/JAFFE/FER-2013/SFEW 2.0/RAF-DB 99.80%/92.92%/54.29%/87.22% CNN 10 fold + Hold out
Zheng et al. (Zheng et al., 2023) 2023 FA RES+Trans+Atten RAF-DB/AffectNet/FERPlus 92.05%/67.31%/90.83% CNN Hold out
Xue et al. (Xue et al., 2021) 2021 FA+FD RES+Trans+Atten RAF-DB/AffectNet/FERPlus 90.91%/66.23%/90.83% CNN Hold out
Liu et al. (Liu et al., 2023a) 2023 FD+FA RES+Trans+Atten RAF-DB/AffectNet/FERPlus 88.21%/60.68%/88.72% CNN Hold out
Meng et al. (Meng and Shi, 2024) 2024 N/A RES+Trans+Atten RAF-DB/AffectNet/AffectNet-8 92.36%/67.44%/64.26% 2sCNN Hold out
Sun et al. (Sun et al., 2023a) 2023 DA RES+Trans+Atten RAF-DB/AffectNet/FER2013 89.50%/65.66%/74.84% 2sCNN Hold out
Feng et al. (Feng et al., 2023) 2023 DA Trans RAF-DB/AffectNet/FERPlus 90.38%/63.33%/90.41% Trans Hold out
Self-supervised Learning Roy et al. (Roy and Etemad, 2021) 2021 DA RES KDEF/DDCF 94.64%/95.26% CNN 10 fold
Roy et al. (Roy and Etemad, 2023) 2023 DA RES KDEF/DDCF/BU3DFE 97.15%/97.34%/97.02% CNN 10 fold
Li et al. (Li et al., 2022c) 2021 FD+DA RES+Atten RAF-DB/CK+/MMI/JAFFE 88.23%/98.77%/79.42%/91.89% CNN 10 fold+Hold out
Wang et al. (Wang et al., 2022a) 2022 DA RES VGGFace2/RAF-DB/FED-RO 60.20%/85.95%/70.00% CNN 10 fold+Hold out
Chen et al. (Chen et al., 2023b) 2023 DA Trans FER2013/AffectNet/SFEW 2.0/RAF-DB 74.95%/66.04%/63.69%/90.98% Trans Hold out
Fang et al. (Fang et al., 2023) 2023 DA RES RAF-DB/FERPlus/AffectNet 78.17%/65.54%/50.33% CNN Hold out
An et al. (An and Jia, 2024) 2024 N/A Trans RAF-DB/FERPlus/AffectNet/AffectNet-8 91.45%/90.16%/63.49%/60.75% Trans Hold out
  • 1DA: Data augmentation; FD: Face detection; CE: Contrast enhancement; IN: Illumination normalization; FA: Face alignment; N/A: Not applicable;
  • 2Trans: Transformer; Atten: Attention; nsCNN: n-stream convolutional neural network; LOSO: Leave-one-subject-out; ROI: Region of interest; RES: Residual; GRU: Gated recurrent unit;
4. Deep MaE Recognition
The ever-increasing data volume has motivated various data-driven methods to handle the technical challenges in computer vision, such as, emotion recognition (Li and Deng, 2020), event detection (Hu et al., 2022) and action detection (Vahdani and Tian, 2022). As an important research direction in computer vision, deep learning-based MaE recognition (a.k.a. deep MaE recognition) has seen its significant advancements in both practical applications and research experiments. More specifically, deep MaE recognition has progressed from analyzing multiple static images to handling complex dynamic image sequences. Hereinafter of this section, we provide an in-depth exploration of deep learning-based MaE recognition methods that are tailored for static image analysis (i.e., deep static MaE recognition) and for dynamic sequential image analysis (i.e., deep dynamic MaE recognition). By examining the state-of-the-art methodologies, we aim to highlight the strengths and advancements in MaE recognition and to emphasize their potential to address increasingly complex wild scenarios.
4.1. Deep Static MaE Recognition
As shown in Table 3, MaE recognition from multiple static images has been proposed due to the efficiency of data processing and the potential for more detailed analysis.
4.1.1. Ensemble learning
Ensemble learning in deep static MaE recognition can exploit complementary information from multiple feature representations of a single emotion image. More specifically, ensemble learning can be used at different stages of the deep static MaE recognition, such as, data preprocessing (Kim et al., 2016), input enhancement (Sharifnejad et al., 2021; Hariri and Farah, 2021; Mohan et al., 2021; Shao and Qian, 2019; Viswanatha Reddy et al., 2020), network generation (Kim et al., 2015; Pons and Masip, 2018b), feature extraction (Xie and Hu, 2019; Wadhawan and Gandhi, 2023), and emotion classification (Mohan et al., 2021; Hua et al., 2019). More specifically, deformation and normalization have been used to preprocess the data before training the deep models for MaE recognition (Kim et al., 2016). When applied at the stage of input enhancement, ensemble learning can improve the MaE recognition by integrating multiple types of textural features, such as, local binary patterns (LBPs) (Sharifnejad et al., 2021; Hariri and Farah, 2021), facial landmark point (Mohan et al., 2021), covariance feature (Shao and Qian, 2019), gradient images (Hariri and Farah, 2021), and gravitational force (Viswanatha Reddy et al., 2020). At the stage of network generation, deep models with various kernel shapes and parameter initializations are ensembled to improve performance (Kim et al., 2015; Pons and Masip, 2018b). For example, Saurav et al. (Saurav et al., 2022) proposed an integrated convolutional neural network (CNN) architecture that leveraged two deep models with distinct kernel shapes. The aggregation of features extracted from deep models represents another prevalent research direction for constructing model ensembles in deep static MaE recognition. Several studies (Xie and Hu, 2019; Wadhawan and Gandhi, 2023) have explored integrating different regions (e.g., eyes, mouth, nose, and the entire image) of the face for model ensembling. Through the different ensemble strategies (Xie and Hu, 2019; Wadhawan and Gandhi, 2023), an integrated feature that combines information from all facial regions can be generated for MaE recognition. In addition to the common integration rules (e.g., majority voting, simple average, and weighted average), the weighted average for decision ensemble was also investigated in (Mohan et al., 2021; Hua et al., 2019). Despite the aforementioned merits, ensemble learning increases the training cost of the model and lacks transparency and interoperability.
4.1.2. Transfer learning
Transfer learning in deep static MaE recognition enables deep models to be fine-tuned on MaE datasets after being pre-trained on large-scale, high-quality MaE datasets. The utilization of transfer learning can alleviate the overfitting issue due to the limited training data for static MaE recognition. By pre-training the model with additional data from other relative tasks, deep static MaE recognition methods can obtain more generic knowledge. For example, Ng et al. (Ng et al., 2015) used ImageNet (Deng et al., 2009) dataset to obtain the pre-trained models that are fine-tuned by the other related MaE datasets. The training efficiency can be improved by only fine-tuning the dense layers of deep models; therefore, the training computing expenditure is reduced in (Akhand et al., 2021). Instead of using general-purpose datasets, MaE-specific datasets can be used for deep model pre-training (Ngo and Yoon, 2020; Atabansi et al., 2021). For instance, Ngo et al. (Ngo and Yoon, 2020) employed the face identification dataset (e.g., VGGFace2 (Cao et al., 2018)) for model transfer. To further mitigate the overfitting in deep static MaE recognition, Atabansi et al. (Atabansi et al., 2021) used another MaE recognition dataset with high resolution and huge quantity for model pre-training.
4.1.3. Multi-task learning
Multi-task learning allows a deep model to extract macro-expression features from static facial images by using other facial behavior analysis tasks as auxiliary tasks in the deep static MaE recognition. For example, recent deep static MaE recognition methods have integrated facial landmark localization (Chen et al., 2021) and AU detection (Pons and Masip, 2018a) to extract more robust MaE features (Pons and Masip, 2018a; Chen et al., 2021; Yu and Xu, 2022; Liu et al., 2024). Moreover, Pons and Masip found that leveraging shared features across AU and MaE can enhance the performance of recognition (Pons and Masip, 2018a). To further improve the synergy, Li et al. (Li et al., 2023) introduced an alignment loss to constrain the feature distribution between the AU detection and MaE recognition tasks. Furthermore, the multi-task learning can be used to fuse the global and local facial expressions by automatically assigning weights based on the importance of global and local facial information (Yu et al., 2020a; Xiao et al., 2023). Besides, facial expression synthesis (Zhao et al., 2021b), head pose estimation (Chen et al., 2022b), body gesture (Zaghbani and Bouhlel, 2022), and gender learning (Pan et al., 2022) had been demonstrated as promising collaborative auxiliary tasks for deep static MaE recognition and could significantly improve performance. Therefore, Chen et al. (Chen et al., 2022b) proposed a dual-attention based multi-task learning framework that contains a separate channel attention mechanism to calculate task-specific attention weights and an orthogonal channel attention loss to optimize the selection of feature channels for each auxiliary task. Since human emotion is conveyed equally via the body and the face in most cases, Zaghbani et al. (Zaghbani and Bouhlel, 2022) combined the upper body gesture action detection task for the deep static MaE recognition. Pan et al. (Pan et al., 2022) incorporated gender learning as an auxiliary task to factor in the effects of gender on the deep static MaE recognition since the characteristics of the same facial expressions from males, females, and infants can be significantly different. Instead of using a single auxiliary task, several studies proposed various multiple tasks for deep static MaE recognition to achieve more synergy. For instance, Qin et al. (Qin et al., 2024) developed an algorithm that simultaneously performs face recognition, MaE recognition, age estimation, and face attribute prediction. Similarly, Foggia et al. (Foggia et al., 2023) proposed a comprehensive multi-task framework to integrate gender, age, ethnicity, and MaE recognition using facial images.
4.1.4. Attention-based learning
Attention-based learning is used to enrich the spatial information in deep static MaE recognition. The attention modules can focus on both global and local regions and generate more comprehensive static facial expression features that typically include local information, global information, and the corresponding dependency. Drawing inspiration from the human ability to locate salient objects in complex visual environments, attention mechanisms have emerged as a powerful tool in deep static MaE recognition (Itti and Koch, 2001). By prioritizing critical facial regions, attention modules can significantly enhance the extraction and integration of expression features in order to achieve a more accurate and robust recognition model. For example, Fernandez et al. (Marrero Fernandez et al., 2019) developed an attention module to enhance the extraction of expression features by assigning higher weights to relevant regions. Li et al. (Li et al., 2020a) integrated LBP and deep features with an attention mechanism to enhance the performance of deep static MaE recognition. In wild scenarios, pose variations and occlusions present significant challenges for deep static MaE recognition (Li et al., 2019b; Wang et al., 2020b). To address the aforementioned issues, various attention modules have been proposed to enhance model robustness and performance. For example, Li et al. (Li et al., 2019b) introduced an attention module designed to identify local facial regions associated with MaEs while simultaneously incorporating complementary global facial information. This dual-focus approach enables more robust MaE recognition under occlusion conditions by balancing localized details with a holistic view of the face. Wang et al. (Wang et al., 2020b) proposed a region attention module that is tailored to address occlusions and pose variations. The region attention module in (Wang et al., 2020b) highlights significant facial regions by leveraging attention weights and combines local region features to generate a compact and fixed-length feature representation for final classification. Motivated by the idea of integrating global and local features, Zhao et al. (Zhao et al., 2021a) developed a global multi-scale and local attention module. Zhao et al. (Zhao et al., 2021a) confirmed that the proposed attention module can ensure a comprehensive representation of facial expressions by capturing global features through multiple receptive fields while guiding the model to focus on salient local features. Motivated by facial attributes and human perception mechanisms, Liu et al. (Liu et al., 2022a) introduced an adaptive multilayer perceptual attention module. By effectively capturing critical information from global, local, and salient facial regions, the proposed module in (Liu et al., 2022a) significantly improves the robustness of deep static MaE recognition.
First proposed in (Dosovitskiy et al., 2020), vision transformers (ViTs) have brought remarkable advancements in the domain of deep static MaE recognition, primarily due to their ability to capture long-range dependencies across image patches (Aouayeb et al., 2021; Zheng et al., 2023). By leveraging the self-attention module inherent in transformers, Aouayeb et al. (Aouayeb et al., 2021) pioneered the integration of ViTs with a squeeze-and-excitation block that can extract expression-related features. Xue et al. (Xue et al., 2021) exploited the ViT architecture to explore relationships among various facial regions for a more holistic understanding of static MaEs. To address the occlusion issue, Liu et al. (Liu et al., 2023a) introduced a patch attention module that assigns attention weights to local facial regions. The integration of the patch attention module with a ViT can effectively capture both local and global dependencies. Building upon the advancements of ViTs, Zheng et al. (Zheng et al., 2023) developed a dual-stream model that leverages a cross-fusion mechanism to combine facial landmark features with image-based features. Furthermore, the incorporation of additional modalities has proven instrumental in augmenting transformer-based deep static MaE recognition. For example, gradient images (Meng and Shi, 2024) and gray-level co-occurrence matrices (Sun et al., 2023a) have been used as complementary inputs to enrich the feature space. These modalities not only enhance the discriminative power of the models but also provide deeper insights into the complex interplay of facial features. The optimization of transformer architectures also plays a critical role in advancing deep static MaE recognition. In (Feng et al., 2023), Feng et al. demonstrated that adjusting the parameter structures of ViTs through optimization algorithms has led to significant performance improvements.
4.1.5. Self-supervised learning
Self-supervised learning can enhance deep static MaE recognition models by broadening their ability to understand and analyze MaEs from a spatial perspective. More specifically, self-supervised learning is utilized to extract meaningful features from unlabeled data that come from unannotated MaEs (Li et al., 2022c; Wang et al., 2022a; Chen et al., 2023b; Fang et al., 2023; An and Jia, 2024), data captured from different perspectives (Roy and Etemad, 2021, 2023; Zbontar et al., 2021), and multi-modal data sources (Siriwardhana et al., 2020; Chaudhari et al., 2023; Halawa et al., 2024).
When applying for unlabeled MaE datasets, Li et al. (Li et al., 2022c) proposed a self-supervised learning method to facilitate compound MaE recognition with multiple labels. To avoid the expenditure of manual annotation, Wang et al. (Wang et al., 2022a) introduced an automatic occluded MaE recognition method that can effectively use large volumes of unlabeled data. Chen et al. (Chen et al., 2023b) combined self-supervised learning with few-shot strategies to train deep models for static MaE recognition with a limited amount of labeled data. Fang et al. (Fang et al., 2023) introduced an innovative approach by leveraging contrastive clustering in self-supervised models. The proposed method in (Fang et al., 2023) can strategically refine pseudo labels within face recognition datasets and thereby unlock the potential to enhance deep static MaE recognition. An et al. (An and Jia, 2024) proposed a self-supervised static MaE recognition that learns multi-level facial features without requiring labeled data.
The self-supervised learning can be employed for deep static MaE recognition tasks using non-frontal data samples. For example, Roy et al. (Roy and Etemad, 2021) proposed a contrastive learning method for multi-view MaEs to address the viewpoint sensitivity and limited quantities of labeled data. By aligning features of the same expression from different perspectives, the proposed contrastive learning method can generate view-invariant embedding features for multi-view static MaE recognition. Later, Roy et al. (Roy and Etemad, 2023) introduced an improved version of the proposed contrastive learning method in (Roy and Etemad, 2021). After obtaining effective view-invariant features, the proposed approach in (Roy and Etemad, 2023) can incorporate supervised contrastive loss and Barlow Twins loss (Zbontar et al., 2021) to further differentiate MaE features with minimized redundancy.
The self-supervised learning had also been applied to multi-modal MaE recognition to address the challenges of integrating data from various modalities without explicit annotation. Siriwardhana et al. (Siriwardhana et al., 2020) utilized pre-trained self-supervised models to extract text, speech, and vision features to improve deep static MaE recognition tasks. Moreover, multi-modal self-supervised learning frameworks had been explored to extract and fuse multi-modal features to achieve promising results in deep static MaE recognition (Chaudhari et al., 2023; Halawa et al., 2024).
4.1.6. Potential limitations of different methods on deep static MaE recognition.
While various methods in deep static MaE recognition have significantly improved model performance and broadened their applicability, they still present certain limitations. Ensemble learning, which combines multiple models to boost performance, often increases computational costs and reduces transparency and interoperability due to added complexity. Transfer learning, which relies on pre-trained models, can be less effective if there is a significant domain gap between the source and target datasets. Moreover, improper fine-tuning can reduce the model’s ability to adapt to new datasets. Multi-task learning enhances model performance by jointly learning MaE recognition alongside related tasks. However, the choice of auxiliary tasks is critical since irrelevant or conflicting tasks may hinder learning. Attention mechanisms, while powerful, can focus excessively on certain regions, potentially ignoring other important information, especially in cases of facial occlusions or pose variations. Self-supervised learning is very dependent on the quality of input data or the limited labeled data. Most current large-scale MaE datasets are collected from different viewing angles, lighting conditions, and occlusions in network environments that cannot be controlled. Therefore, if the unlabeled data contains much noise or atypical samples, then the model may learn incorrect feature representations, which impairs its final recognition performance.
Table 4. An Overview of Representative Methods on Deep Dynamic MaE Recognition
Learning Paradigms Method Year Pre-processing Block Dataset Performance
Network
Structure
Protocal
Ensemble Learning Kahou et al. (Kahou et al., 2013) 2013 FA+DA N/A AFEW2 41.03% CNN Cross validation
Kahou et al. (Kahou et al., 2016) 2016 FA+DA N/A AFEW4 47.67% CNN 2 fold
Liu et al. (Liu et al., 2022b) 2022 DA RES+Atten BU-3DFE/MMI/AFEW/DFEW 85.33%/91%/53.98%/65.35% CNN Cross validation
Liu et al. (Liu et al., 2023b) 2023 DA Trans BU-3DFE/MMI/AFEW/DFEW 88.17%/92.5%/54.26%/65.85% CNN Cross validation
Pan et al. (Pan et al., 2024) 2024 FA RES+Atten eNTERFACE’05/BAUM-1s/AFEW 54.6%/57.5%/54.7% CNN Cross validation
Explicit Spatio-temporal Learning Zhang et al. (Zhang et al., 2017) 2017 DA N/A CK+/Oulu-CASIA/MMI 45.5%/25.8%/24.4% CNN+RNN 10 fold
Zhi et al. (Zhi and Wan, 2019) 2019 DA N/A Biovid Heat Pain/CASME II/SMIC 29.7%/54.6%/57.8% RNN LOGO
Lee et al. (Lee et al., 2020) 2020 DA Atten MAVFER RMSE/CC/CCC:0.112,0.563,0.521 RNN+CNN Hold out
Hasani et al. (Hasani and Mahoor, 2017) 2017 FA+DA RES CK+/MMI/FERA/DISFA 67.52%/54.76%/41.93%/40.51% CNN+RNN 5 fold
Deng et al. (Deng et al., 2019) 2019 N/A RES CK+/MMI/AFEW 94.39%/80.43%/82.36% CNN+RNN 5 fold
Khanna et al. (Khanna et al., 2024) 2024 DA RES Ravdess/CK+/Baum1 91.69%/98.61%/73.73% CNN Hold out
Multi-task Learning Yu et al. (Yu et al., 2020b) 2020 FA+FD N/A CK+/Oulu-CASIA 98.77/90.40% CNN+LSTM 10 fold
Jin et al. (Jin et al., 2021) 2021 DA RES Aff-wild2 89.87% CNN Hold out
Xie et al. (Xie et al., 2023) 2023 FA+DA+ROI Trans+Atten MU-MOSI/CMU-MOSEI 84.61%/84.61% BERT Hold out
Attention-Based Learning Meng et al. (Meng et al., 2019) 2019 FD+FA RES+Atten CK+/AFEW 8.0 99.69%/51.8% CNN 10 fold
Liu et al. (Liu et al., 2020a) 2020 FD+DA Atten CK+/Oulu-CASIA/MMI/AffectNet 99.54%/88.33%/87.06%/63.71% CNN+RNN 10 fold
Sun et al. (Sun et al., 2021) 2021 FD Atten CK+/MMI /Oulu-CASIA 99.1%/89.88%/87.33% 2sCNN 10 fold
Xia et al. (Xia et al., 2022) 2022 FD Atten Aff-Wild2/RML/AFEW 50.3%/78.32%/59.79% CNN+3DCNN Hold out + LOSO
Chen et al. (Chen et al., 2023a) 2020 FD+DA Atten CK+/Oulu-CASIA/MMI 99.08%/91.25%/82.21% 3D CNN 10 fold
Zhao et al. (Zhao and Liu, 2021) 2021 FD+DA RES+Atten DFEW/AFEW
UAR: 53.69%/47.42
WAR:65.70%/50.92
CNN 5 fold+Hold out
Huang, et al. (Huang et al., 2021) 2021 FA RES+Atten CK+/FERplus/RAF-DB 100%/90.04%/88.26% CNN 10 fold+Hold out
Ma et al. (Ma et al., 2022) 2022 FD+FA RES+Trans DFEW/AFEW
UAR: 54.58%/49.11%
WAR: 66.65%/54.23%
CNN 5 fold+Hold out
Zhang et al. (Zhang et al., 2022) 2022 FD RES+Trans ABAW F1: 35.9% CNN Hold out
Zhao et al. (Zhao et al., 2023b) 2022 FA+FD Atten+Trans
CK+/Oulu-CASIA/eNTERFACE05/
AFEW/CAER
98.78%/89.17%/54.62%/
51.17%/77.06%
GCN Cross validation
Wang et al. (Wang et al., 2024) 2024 FA RES+Trans DFEW/FERV39k
UAR: 60.28%/41.28%
WAR: 71.42%/51.02%
CNN N/A
Chen et al. (Chen et al., 2024) 2024 DA RES+Trans DFEW/FERV39k/AFEW 8.0
UAR: 58.65%/41.91%/52.23%
WAR: 69.91%/50.76%/55.40%
CNN Hold out+5 fold
Zhang et al. (Zhang et al., 2022) 2023 FD+DA Trans MAFW11/MAFW43/DFEW/AFEW
UAR: 39.37%/15.22%/57.16%/50.22%
WAR: 52.85%/39.00%/68.85% /52.96%
CNN+Trans 5 fold+Hold out
Self-supervised Learning Song et al. (Song et al., 2021) 2021 FD+ROI N/A SEMAINE MSE:0.058/0.072 UNet Cross validation
Sun et al. (Sun et al., 2023b) 2023 ROI Trans DFEW/FERV39k/MAFW
UAR:63.41%/43.12%/41.62%
WAR:74.43%/52.07%/54.31%
Trans Cross validation
Chumachenko et al. (Chumachenko et al., 2024) 2024 FD Trans DFEW/MAFW
UAR:66.85%/44.25%
WAR:77.43%/58.45%
Trans 5 fold
  • 1DA: Data augmentation; FD: Face detection; CE: Contrast enhancement; FA: Face alignment; N/A: Not applicable;
  • 2Trans: Transformer; Atten: Attention; nsCNN: n-stream CNN; LOSO: Leave-one-subject-out; RNN: Recurrent neural network; LSTM: Long short-term memory; BERT: Bidirectional encoder representations from transformer; GCN: Graph convolutional network;
  • 3 UAR: Mean class accuracy; WAR: Overall accuracy; RMSE: Root mean squared error; CC: Correlation coefficient; CCC: Concordance correlation coefficient;
4.2. Deep Dynamic MaE Recognition
While MaE recognition can be achieved using static images and isolated frames, it greatly benefits from analyzing consecutive dynamic frames over time. In this section, we provide a comprehensive review of recent deep dynamic MaE recognition methods, summarized in Table 4.
4.2.1. Ensemble learning
Ensemble learning can enhance the accuracy and robustness of dynamic MaE recognition by frame aggregation that can combine temporal dynamic features across multiple frames and frame-level emotion classification results In deep dynamic MaE recognition, the frame aggregation can be implemented at decision-level (Kahou et al., 2013, 2016) and feature-level (Nguyen et al., 2022; Liu et al., 2022b, 2023b; Pan et al., 2024).
The decision-level frame aggregation integrates the classification results of individual frames in the form of class probability vectors. For instance, Kahou et al. (Kahou et al., 2013) explored the decision-level frame aggregation by using averaging and expansion for deep dynamic MaE recognition. Later, Kahou et al. (Kahou et al., 2016) used expansion or contraction methods to aggregate single-frame probabilities into fixed-length video descriptors. Since the number of frames may vary, statistical characteristics (e.g., mean, variance, minimum, and maximum) can be utilized to summarize the frame-level outputs. However, relying solely on per-frame decisions may overlook the temporal dependencies between consecutive frames, which may impact the performance of deep dynamic MaE recognition.
The feature-level frame aggregation focuses on integrating frame-level features for final prediction. Nguyen et al. (Nguyen et al., 2022) concatenated frame features and fed them into a 3D CNN model for MaE recognition. By leveraging the attention mechanism, Liu et al. (Liu et al., 2022b) introduced a frame aggregation method to integrate expression-related features. Specifically, they segmented videos into short clips and employed an attention-based feature extractor to capture salient features from these segments. Then, an emotional intensity activation network was designed to locate salient clips for generating robust features. Building on this, Liu et al. (Liu et al., 2023b) proposed a transformer-based frame aggregation method for dynamic MaE recognition. To effectively integrate interrelationships among multi-cue features, Pan et al. (Pan et al., 2024) developed a hybrid fusion method for video-based MaE recognition. The proposed ensemble method in (Pan et al., 2024) combines the strengths of different types of features (e.g., appearance, geometry, and high-level semantic knowledge) and thereby improves the overall performance and robustness of dynamic MaE recognition systems.
4.2.2. Explicit spatio-temporal learning
Explicit spatio-temporal learning focuses on constructing deep models that are designed to extract spatio-temporal information from the dynamic MaE datasets. The explicit spatio-temporal learning methods can incorporate the temporal information into the encoded features by using a sequence of frames within a sliding window. The recurrent neural network (RNN) (Elman, 1990) and its advanced version (i.e., long short-term memory, LSTM) (Hochreiter and Schmidhuber, 1997) have been used in dynamic MaE recognition due to the ability to process sequential data (Zhang et al., 2017; Lee et al., 2020; Wang et al., 2022b; Uddin et al., 2021; ZhouXuan, 2021; Pei et al., 2021; Yu et al., 2018b; An and Liu, 2020; Abdullah et al., 2020). For instance, Zhang et al. (Zhang et al., 2017) employed an RNN to articulate morphological changes and dynamic evolution of MaEs by exploiting key facial regions based on facial landmarks. Zhi et al. (Zhi and Wan, 2019) proposed a lightweight LSTM-based method with less complexity to produce a sparse representation for dynamic MaE recognition. Lee et al. (Lee et al., 2020) introduced attention-guided convolutional LSTM to capture dynamic spatio-temporal information. By leveraging the depth and thermal sequences as guidance priors, the proposed model can guide the model to focus on discriminative visual regions. An LSTM autoencoder was used to learn temporal dynamics for dynamic MaE recognition (Uddin et al., 2021). Besides, the 3D deep models (e.g., 3D CNN (Ji et al., 2012)) have also been adopted to extract spatio-temporal features directly for dynamic MaE recognition (Hasani and Mahoor, 2017; Deng et al., 2019; Cai et al., 2018). Hasani et al. (Hasani and Mahoor, 2017) combined 3D Inception-ResNets (Szegedy et al., 2017) with an LSTM unit to capture spatio-temporal information. Deng et al. (Deng et al., 2019) proposed a 3D CNN framework consisting of a stem layer, 3D Inception-ResNets structure, and gated recurrent unit (GRU) layer to process video data. To further improve the performance of the model, the island loss (Cai et al., 2018) was incorporated to increase inter-class differences while minimizing intra-class variations. Khanna et al. (Khanna et al., 2024) developed an end-to-end spatio-temporal deep model that integrates residual networks (He et al., 2016) and DenseNet (Huang et al., 2017) for MaE recognition.
4.2.3. Multi-task learning
Multi-task learning in deep dynamic MaE recognition focuses on capturing the dynamic variation patterns of MaEs over time in order to enhance the understanding of expression changes in video clips. By fusing local dynamic features extracted from a part-based module, Yu et al. (Yu et al., 2020b) proposed a multi-task module that can capture the subtle variations of MaEs from both local and global features. Jin et al. (Jin et al., 2021) utilized a multi-task framework involving AU detection and MaE recognition to pre-train the visual feature extractor. Then, both visual and audio features from videos were concatenated and fed into a transformer encoder to extract temporal features for recognition. Xie et al. (Xie et al., 2023) expanded the scope of auxiliary tasks by incorporating dynamic MaE recognition from multiple modalities, such as text, audio, and video.
4.2.4. Attention-based learning
Attention-based learning in deep dynamic MaE recognition focuses on both spatial and temporal information. Compared with attention modules in deep static MaE recognition, attention modules in deep dynamic MaE recognition operate across the temporal and even channel dimensions to extract richer video features. Furthermore, these attention modules can integrate multimodal features to derive more comprehensive spatio-temporal representations.
Various attention modules have been proposed to enhance deep dynamic MaE recognition by integrating spatial, temporal, and channel-wise information, e.g., self-attention (Meng et al., 2019), relation attention (Meng et al., 2019), AU attention (Liu et al., 2020a), and multi-attention (Sun et al., 2021; Xia et al., 2022; Chen et al., 2023a). For example, Meng et al. (Meng et al., 2019) proposed a self-attention module to aggregate all input frame features and a relation attention module to capture informative features from global and local contexts. Liu et al. (Liu et al., 2020a) introduced an AU attention mechanism to augment long-range expression information by focusing on specific AU regions. Besides, multi-attention is implemented in (Sun et al., 2021; Xia et al., 2022; Chen et al., 2023a) to obtain the rich complementary information in deep dynamic MaE recognition. For instance, Sun et al. (Sun et al., 2021) proposed two attention modules to integrate facial features extracted from two deep models. Subsequently, an additional attention-based deep model was introduced to compose a comprehensive feature representation from MaE sequences. Xia et al. (Xia et al., 2022) developed a multi-attention module for dynamic MaE recognition. The proposed module in (Xia et al., 2022) can integrate spatial and temporal information from areas of interest (e.g., expressive frames and salient facial patches). Chen et al. (Chen et al., 2023a) further expanded the scope by incorporating channel-wise attention and demonstrated its effectiveness in enhancing both performance and interpretability. Specifically, the proposed method in (Chen et al., 2023a) employed a spatio-temporal module and a channel attention module to explore correlations across spatio-temporal and channel dimensions. Moreover, the proposed method in (Chen et al., 2023a) can generate spatio-temporal attention maps for visualization.
Transformers have been used to enhance dynamic MaE recognition by integrating spatial, temporal, and contextual video features (Zhao and Liu, 2021; Huang et al., 2021; Ma et al., 2022; Zhang et al., 2022). Zhao et al. (Zhao and Liu, 2021) proposed a dual-transformer framework that extracts robust MaE features by leveraging a spatial transformer module to capture spatial configurations and a temporal transformer module to account for temporal changes in expression videos. Huang et al. (Huang et al., 2021) developed a MaE recognition framework that utilizes grid-wise attention to capture dependencies among facial regions and a ViT module to extract global features. To address long-range dependencies within videos, Ma et al. (Ma et al., 2022) designed a spatio-temporal transformer that integrates contextual relationships across frames, providing a comprehensive representation of dynamic MaE features. Additionally, some studies (Zhao et al., 2023b; Wang et al., 2024; Chen et al., 2024) incorporated external knowledge into transformers to enhance feature extraction. For instance, Zhao et al. (Zhao et al., 2023b) employed facial landmarks to construct spatio-temporal graphs, using graph convolutional blocks and transformer modules to extract MaE-related features. Wang et al. (Wang et al., 2024) combined multi-scale spatial information from CNNs with temporal features extracted from transformer modules. Similarly, Chen et al. (Chen et al., 2024) introduced multi-geometry knowledge into transformers, enabling the extraction of spatio-temporal geometry information for MaE recognition.
Moreover, transformers have been leveraged to integrate multi-modal features for dynamic MaE recognition. For instance, Zhang et al. (Zhang et al., 2022) proposed a transformer-based framework that fuses static visual features with dynamic multi-modal information derived from text, audio, and video. In the proposed framework of (Zhang et al., 2022), static frame features act as a guiding mechanism for the multi-modal fusion process to ensure the alignment of temporal dynamics with spatial cues from individual frames. Building on the framework in (Zhang et al., 2022), Zhang et al. (Zhang et al., 2024) introduced an advanced multi-modal transformer framework with improved methods for feature fusion and extraction. Specifically, the framework employs three transformer-based encoders, each dedicated to extracting features from a specific modality—visual, audio, or text—while enabling modality-specific adaptation. Unlike the earlier framework (Zhang et al., 2022), which focused on integrating multi-modal data using static frame guidance, the subsequent method (Zhang et al., 2024) achieves more robust multi-modal feature integration by aligning semantic information across modalities. This is accomplished by mapping features from all modalities into a shared latent space anchored in visual features, thereby supporting a more unified representation of multi-modal data.
4.2.5. Self-supervised learning
Self-supervised learning focuses on capturing changes in expressions across spatial and temporal dimensions in the deep dynamic MaE recognition. Moreover, self-supervised models can learn spatio-temporal features from unlabeled video data. Since the self-supervised learning models can interpret videos by learning the order of frames in human action recognition (Fernando et al., 2017; Sharma et al., 2019), Song et al. (Song et al., 2021) introduced a self-supervised framework with a rank loss mechanism to generate dynamic MaE features. By sorting preceding and following frames based on their distance from a central frame, the proposed framework can generate robust dynamic MaE features. Sun et al. (Sun et al., 2023b) proposed a large-scale self-supervised model for pre-training a dynamic MaE recognition system using extensive unlabeled facial video data. Chumachenko et al. (Chumachenko et al., 2024) proposed a multi-modal approach that combined complementary features from diverse modalities. The multi-modal approach employed two self-supervised learning encoders pre-trained on audio sequences and static images, and then fine-tuned using audio-visual dynamic videos.
4.2.6. Potential limitations of different methods on deep dynamic MaE recognition.
Ensemble learning in dynamic MaE recognition increases both computational cost and system complexity. For example, feature-level frame aggregation requires substantial computation due to the large volume of frame-level features and complex network structures e.g., 3D CNNs, attention mechanisms, or transformers. Explicit spatiotemporal learning methods, such as 3D CNNs and RNN variants such as LSTM, GRU, are highly computationally expensive during training and inference, which may be impractical for resource-constrained applications. Furthermore, these models for the explicit spatiotemporal learning method with large amounts of parameters, are prone to overfitting during the training stage, especially when the MaE data is limited and the MaE changes are subtle within video sequences. The selection of appropriate auxiliary tasks in multi-task learning is a major challenge for MaE recognition. Irrelevant tasks or auxiliary tasks with significantly different levels of difficulty can disrupt model training. In multi-modal and multi-task dynamic recognition models, integrating data from text, audio, and video requires precise alignment and synchronization. Misalignment between these modalities can degrade the model’s performance, particularly when capturing subtle expression changes. Attention mechanisms can effectively integrate facial region information and capture long-range dependences, but they can also introduce latency issues, particularly in models like transformers, which suffer from increased computational and memory requirements as the input sequence length grows. Although self-supervised learning captures sequential information, it struggles to model long-term correlations, especially when MaE changes are not immediately obvious.
5. Holographic MiE Analysis
Detecting MiEs is challenging due to their short duration and subtle emotional cues; therefore, each MiE must first be identified before proceeding to recognition. As a foundational step, MiE spotting involves identifying the temporal and spatial occurrences of these brief and subtle MiEs within a video clip. Building on MiE spotting, MiE recognition then analyzes and classifies the detected MiEs. In the remainder of this section, we detail the state-of-the-art approaches to the two major steps of holographic MiE analysis, i.e. MiE spotting and MiE recognition.
5.1. MiE Spotting
MiE spotting, referring to locating temporal segments in a specific video clip, is an important step for holographic MiE analysis. The key description of MiE in a video clip is to identify three time spots: (1) the onset refers to the initial frame where the emotion of an MiE becomes discernible; (2) the apex frame is the frame within the MiE clip that exhibits the highest intensity of emotion; and (3) the offset frame marks the end of the emotion. The current works on MiE spotting can be categorized into apex-frame spotting and interval spotting.
5.1.1. Apex frame spotting
The apex frame reflects the highest intensity of an MiE within a video clip and implies the ground-truth emotion of individuals. Traditional apex frame spotting methods exploited the algorithms based on handcrafted features, such as the descriptors of LBP and its variants due to their effectiveness and robustness. Pfister et al. (Pfister et al., 2011b) pioneered a MiE spotting approach for short spontaneous expressions. The proposed method in (Pfister et al., 2011b) leverages local binary patterns from three orthogonal planes (LBP-TOP) (Zhao and Pietikainen, 2007) to achieve notable results compared to manual detection methods. Yan et al. (Yan et al., 2015) proposed to utilize two handcrafted descriptors (i.e., the constrained local model (Cristinacce et al., 2006) and LBP) to detect the ROI of each face and to extract subsequent features for MiE analysis. To address the limitations of LBP-TOP in capturing essential information, Esmaeili et al. (Esmaeili and Shahdi, 2020) proposed Cubic-LBP—an improved version of LBP. Compared with LBP, Cubic-LBP can extract fine-grained features from 15 planes and integrate the obtained histograms. Although the use of additional planes could enrich the features, Cubic-LBP also introduced redundant data that increases computational expenditure. To alleviate the computational expenditure, Esmaeili et al. (Esmaeili et al., 2022) proposed an intelligent Cubic-LBP by leveraging CNNs to select relevant planes.
MiE spotting can be advanced through frequency-domain features (Li et al., 2018, 2020), CNN-based sliding windows (Zhang et al., 2018a), and attention mechanisms (Yee et al., 2022) for enhanced performance and feature selection. Given the challenges in capturing MiE motion within the spatio-temporal domain, an alternative approach is to extract features from the frequency domain. Li et al. (Li et al., 2018, 2020) introduced a spotting method leveraging the frequency domain to identify the apex frame. The proposed methods in (Li et al., 2018, 2020) demonstrated the effectiveness of the frequency domain feature in MiE spotting. The CNN models have also gained traction in MiE spotting. Zhang et al. (Zhang et al., 2018a) first employed CNNs with a sliding window approach to locate apex frames in long videos. To spot apex frames in onset-offset temporal sequences, Yee et al. (Yee et al., 2022) used an attention module to highlight key regions in optical flow. The proposed method in (Yee et al., 2022) underscores that the attention mechanism has the capability of automatic feature selection in respective regions; therefore, the spotting performance can be improved.
5.1.2. Interval spotting
Interval spotting refers to detecting the onset and offset in a long MiE video with various constraints, such as, MaE, head movement, and blinks. While more challenging than apex frame spotting, interval spotting has broader practical applications. For interval spotting, traditional methods typically set thresholds to locate key points by comparing feature differences across entire videos. For example, Li et al. (Li et al., 2017b) proposed a method based on comparison of difference characteristics. Specifically, the proposed method identified LBP as a more effective baseline technique compared to the optical flow histogram. By incorporating the multi-scale analysis and a sliding window, Tran et al. (Tran et al., 2017) scaled the video sequence and obtained corresponding samples for spotting MiE intervals. Optical flow has been used in interval spotting due to its ability to describe motion information. Patel et al. (Patel et al., 2015) introduced a heuristic algorithm with optical flow for MiE interval spotting. The algorithm emphasized the continuity of motion direction. By analyzing the direction of motion vectors over time, it ensured that detected movements extracted by optical flow were consistent.
Wang et al. (Wang et al., 2016, 2017) introduced the main directional maximal difference (MDMD) to characterize facial motion by calculating the largest magnitude difference in the main direction of optical flow. The proposed method was later adopted as a baseline in the third facial MiE grand challenge (Li et al., 2020b). Han et al. (Han et al., 2018) enhanced feature difference analysis by combining LBP with the main directional mean optical flow (MDMO) (Liu et al., 2015), which leveraged both texture and motion features for complementary insights. Guo et al. (Guo et al., 2021) extended this work by incorporating both magnitude and angle information from optical flow to detect local MiE movements.
MiE interval spotting methods focusing on handcrafted features and optical flow mentioned above require appropriate threshold settings, which is susceptible to the selected features and unexpected facial motion. To address the challenges, deep model approaches have emerged as promising alternatives. To distinguish the MaE and MiE in a long video sequence, Pan et al. (Pan et al., 2020) presented a local bilinear CNN to detect the fine-grained facial regions associated with MiE. Takalkar et al. (Takalkar et al., 2021) developed LGAttNet, a dual-attention framework that integrated local and global facial features for improved spotting. Gu et al. (Gu et al., 2023) proposed a lightweight deep model to predict the likelihood of a frame being part of a MiE interval for overfitting mitigation. To capture spatio-temporal features, Wang et al. (Wang et al., 2021) designed a 2D spatial and 1D temporal convolutional model. In addition, many studies (Yu et al., 2021a; Yap et al., 2022) utilized 3D CNNs to extract robust spatio-temporal features. Zhou et al. (Zhou et al., 2022) further enhanced performance by integrating 3D CNNs with bidirectional encoder representations from transformers (BERTs) for onset-offset detection. Moreover, approaches leveraging AU analysis (Yin et al., 2023; Yang et al., 2023) can incorporate prior knowledge and reduce noise interference to improve MiE spotting performance.
5.1.3. Potential limitations of different methods on MiE spotting.
Traditional handcrafted descriptors are robust and offer good interoperability. For apex frame spotting, methods like LBP-TOP can handle low-intensity MiEs effectively. However, handcrafted descriptors are typically limited to local information and may fail to fully capture global dynamic patterns. CNN-based sliding window approaches can locate apex frames effectively, but are less time-efficient for long videos. The performance of sliding window methods is sensitive to the choice of window size and step size, impacting both detection speed and accuracy. Apex frame spotting also requires high-quality data, which may not always be available in practical applications.
Traditional interval spotting methods using handcrafted descriptors rely on threshold-based detection to locate key points by comparing feature differences across entire videos. This makes them highly sensitive to feature selection and unexpected facial movements. The fine-tuning of thresholds is a challenge in diverse real-world scenarios. Deep learning models, although successful in MiE interval spotting, require large, labeled datasets. Given that MiEs are subtle and brief, models can overfit to the specific characteristics of the training data.
5.2. MiE Recognition
Since the seminar work (Pfister et al., 2011a), MiE recognition has seen unprecedented development that ranges from shallow machine learning with handcrafted descriptors to deep learning. Hereinafter of this part, the MiE recognition will be thoroughly reviewed.
5.2.1. Shallow machine learning methods
The success of shallow machine learning [^1] The word “shallow” is used to differentiate from the deep learning methods. demands for the well-designed handcrafted descriptors (e.g., LBPs and grayscale images) to extract local and global facial information. As a key component of MiE recognition, the development of descriptors has attracted several research endeavors, such as, LBP with six intersection points (Wang et al., 2015b) and LBP with three mean orthogonal planes (Wang et al., 2015a). More specifically, Wang et al. (Wang et al., 2015b) proposed the LBP with six intersection points to reduce the redundancy inherent in LBP-TOP. Later, Wang et al. (Wang et al., 2015a) proposed the LBP with three mean orthogonal planes to alleviate the computational complexity issue of LBP-TOP. Compared with the LBP-TOP descriptor, the proposed LBP with three mean orthogonal planes offers faster processing with a competitive performance (Wang et al., 2015a). Huang et al. (Huang et al., 2016) developed the spatio-temporal completed local quantization pattern descriptor that incorporates additional information on sign, magnitude, and orientation components to enhance MiE recognition. To address the limitation of LBP-TOP on capturing muscle movement in the oblique direction, Wei et al. (Wei et al., 2022a) proposed a descriptor named as LBP with five intersecting planes (LBP-FIP). The LBP-FIP descriptor enhances the LBP-TOP with an innovative feature (i.e., eight vertices LBP) in order to enrich the feature information by capturing detailed MiE dynamics in multiple directions. Therefore, the LBP-FIP can provide a richer and more effective representation for MiE recognition than the LBP-TOP descriptor.
Besides the series of LBP descriptors, the histogram of gradient (HOG) emerged as another prominent descriptor for MiE recognition (Dalal and Triggs, 2005). By computing and aggregating the gradient directions in local regions of a facial image, HOG demonstrates its robustness to variations in illumination and deformation. Therefore, various HOG-based descriptors were adapted for MiE recognition (Li et al., 2017b; Niu et al., 2018; Zhang et al., 2020). Li et al. (Li et al., 2017b) proposed the histogram of image gradient orientation, which further suppressed illumination issues by discarding magnitude weighting from the first-order derivative. Niu et al. (Niu et al., 2018) introduced a local second-order gradient pattern to capture subtle facial changes in brief video clips. Zhang et al. (Zhang et al., 2020) leveraged the histograms of oriented gradients on three orthogonal planes and the histograms of image-oriented gradients on three orthogonal planes to construct an effective MiE recognition framework.
Due to the superiority of describing brightness patterns in images, optical flow (Horn and Schunck, 1981) can also be used to capture subtle muscle motion from adjective frames for MiE recognition. More specifically, Liu et al. (Liu et al., 2015) introduced the MDMO feature descriptor for MiE recognition. MDMO feature descriptor can integrate local motion information and spatial location from partitioned ROIs on the face, providing a robust representation of facial movements. Liong et al. (Liong et al., 2014) proposed to use the optical strain, consisting of the shear and normal strain tensors from optical flow, for MiE recognition. Happy et al. (Happy and Routray, 2017, 2018) developed a fuzzy histogram of optical flow orientation to encode the temporal patterns of facial micro-movements; therefore, distinctive features that were sensitive to the nuances of MiE can be extracted. Considering the directional continuity of facial motion, Patel et al. (Patel et al., 2015) proposed an optical flow-based method that extracts features from local spatial regions with spatio-temporal integration.
Besides the feature extraction strategies, several works have investigated feature selection (Happy and Routray, 2018; Wang et al., 2022e) and feature fusion (Yu et al., 2018a; Zhang et al., 2020) for MiE recognition. By selecting a subset of relevant features, feature selection enhances the MiE recognition methods by reducing the redundant high-dimensional features. Happy et al. (Happy and Routray, 2018) utilized a fuzzy histogram of optical flow orientation descriptors to capture temporal patterns associated with facial micro-movements. Then, the proposed method explores various feature selection methods to reduce the dimension of feature space. To extract effective spatio-temporal features, Yu et al. (Yu et al., 2018a) combined the improved local directional number pattern in the spatial domain with the pyramid of histograms of orientation gradients in the temporal domain. Zhang et al. (Zhang et al., 2020) proposed a feature selection method to integrate the effective feature components from the seven local regions of the face. Wang et al. (Wang et al., 2022e) introduced an integral projection method to enhance information density, followed by a fixed-point rotation-based feature selection approach to identify features with significant motion variations. To address redundant or misleading features in recognition tasks, Wei et al. (Wei et al., 2022b) implemented a kernelized two-group sparse learning model to optimize feature discrimination.
In general, shallow machine learning methods are based on handcrafted descriptors that demand for extensive prior knowledge on MiEs. After extracting discriminative features, classifying methods (e.g., support vector machine (Yu et al., 2018a) and $k$-nearest neighbor (Happy and Routray, 2018)) are used for final MiE recognition.
5.2.2. Deep learning methods
Compared with the shallow machine learning methods that demand for handcrafted descriptors to extract features, the deep learning for MiE recognition (a.k.a., deep MiE recognition) can directly learn the advanced semantic features from corresponding data and achieve state-of-the-art performance on MiE recognition. Before proceeding, we first discuss the different types of input for MiE recognition.
  • Static images and dynamic frames. Given the subtle motions in per MiE, the apex with peak emotional intensity is often employed to reduce computational complexity (Zhou et al., 2019; Gan and Liong, 2018; Liong et al., 2018). Li et al. (Li et al., 2018) demonstrated the effectiveness of deep models using apex frames for MiE recognition. Furthermore, Sun et al. (Sun et al., 2020) showed that apex frames performed better than complete video sequences in certain scenarios. To capture dynamic information, several approaches utilize multiple sparse frames (Liu et al., 2020b; Gan et al., 2019; Xia et al., 2020a; Zhu et al., 2023) or dynamic images (Nie et al., 2021; Le et al., 2020; Verma et al., 2019; Sun et al., 2020). Sparse frames allow for the extraction of richer information to describe MiE motion. Liu et al. (Liu et al., 2020b) proposed a dynamic segmented sparse imaging module to highlight key frames and describe subtle motion changes. Gan et al. (Gan et al., 2019) presented OFF-ApexNet to extract optical flow features from the onset to apex frames. In addition, Xia et al. (Xia et al., 2020a) calculated optical flow maps from the onset to apex frames. Zhu et al. (Zhu et al., 2023) proposed Learning to rank onset-occurring-offset features, which used onset, offset, and a random frame to reduce low-intensity information and enhance emotional expressiveness, achieving discriminative features.
  • Temporal information. Inspired by an action recognition method that converted dynamic clips into single dynamic images to integrate spatio-temporal information (Bilen et al., 2016), multiple MiE recognition approaches (Nie et al., 2021; Le et al., 2020; Verma et al., 2019; Sun et al., 2020) use dynamic images to capture subtle movements during expression clips. Due to the incorporation of richer temporal information without incurring significant computational overhead, the methods using dynamic images can achieve better performance than apex frame-based approaches in MiE recognition.
  • Complete clips. Advances in deep models for time series processing allow extraction of spatio-temporal information from continuous clips and several methods demonstrate competitive performance (Li et al., 2019a; Zhao et al., 2021c; Reddy et al., 2019; Bai and Goecke, 2020). However, due to information redundancy and the limited scale of MiE datasets, such deep models may lead to high computational costs and overfitting.
Different deep learning paradigms have been explored to address various challenges in MiE recognition (Fan et al., 2023; Nguyen et al., 2023; Das et al., 2021). Therefore, we are motivated to discuss various learning paradigms for deep MiE recognition.
Transfer learning. Due to the scarcity of MiE datasets and the subtle intensity of MiEs, transfer learning for deep MiE recognition needs a stronger reliance on prior knowledge on facial expressions and the ability to capture subtle motion changes. Transfer learning in MiE recognition tends to learn features from MaEs and AUs. For instance, Patel et al. (Patel et al., 2016) pre-trained models on both ImageNet and MaE datasets and demonstrated that MaE datasets are more suitable for MiE recognition due to their domain-specific features. The application of transfer learning to deep MiE recognition involves two steps: (1) pre-training on related datasets (e.g., facial expressions and ImageNet datasets) to learn general features of human faces and objects; and (2) fine-tuning on MiE datasets (Li et al., 2020; Liu et al., 2020b; Chen et al., 2020b). Moreover, transfer learning has been investigated by leveraging knowledge from MaE datasets that contain a large number of labeled MaEs (Peng et al., 2018; Zhi et al., 2019; Wang et al., 2020a; Xia et al., 2020b; Xia and Wang, 2021). For example, Zhi et al. (Jia et al., 2018) proposed to pre-train a 3D-CNN on the Oulu-CASIA dataset to enrich MiE features and improve recognition performance. Before fine-tuning on MiE datasets, Wang et al. (Wang et al., 2020a) proposed to address the overfitting issue by sequentially pre-training network parameters on ImageNet and MaE datasets. To better utilize dynamic muscle motion, Xia et al. (Xia and Wang, 2021) proposed a dual-stream deep model with one stream pretrained on MiE data and the other on MaE data. The domain discriminator and the relation classifier were designed to account for spatio-temporal information during the transfer process. In particular, the domain discriminator module focuses on capturing the static textures of facial appearances, and the relation classifier predicts the correct relation among temporal features with different sampling intervals from the double streams of the model. Indolia et al. (Indolia et al., 2022) revealed that fine-tuning a pre-trained ResNet18 model with a self-attention mechanism can identify relevant facial regions for MiE recognition. Liu et al.(Liu et al., 2022c) used MaE samples to train the feature extractor and optimized the hyper-parameters through grid search for subsequent deep MiE recognition.
As an alternative of fine-tuning, knowledge distillation involves transferring knowledge from a complex “teacher” model to a simpler “student” model. For example, Sun et al. (Sun et al., 2020) used a pre-trained teacher model to distill AU-based knowledge to guide the student model. Li et al. (Li et al., 2021b) developed a dual-view attentive similarity-preserving distillation approach to learn AU knowledge and employed a semi-supervised co-training method to generalize the teacher model. Song et al. (Song et al., 2022) incorporated related expression samples into a first-order motion model to extract motion variations from neutral expressions to MaE and MiE. More specifically, Song et al. (Song et al., 2022) used a dual-channel encoder-decoder model guided by teacher model features to learn the transition features from MiE to MaE.
Multi-task learning. Multi-task learning in deep MiE recognition focuses on incorporating fine-grained auxiliary tasks to enhance the model’s sensitivity to subtle facial changes. Specifically, multi-task learning leverages auxiliary tasks (e.g., gender detection (Li et al., 2019c; Nie et al., 2021) and AU detection (Fu et al., 2022; Wang et al., 2023b)) to improve model generalization and robustness in deep MiE recognition. Li et al. (Li et al., 2019c) utilized multiple related tasks (e.g., facial landmark detection, gender detection, smiling recognition, profile analysis, and glasses detection) as auxiliary tasks for MiE recognition. Similarly, Nie et al. (Nie et al., 2021) proposed a dual-stream approach based on gender for MiE recognition, demonstrating that integrating gender features can enhance MiE recognition performance. As an auxiliary task in MiE recognition, Fu et al. (Fu et al., 2022) employed AU detection to extract fine-grained features, thereby improving the capabilities of attention-based models. Wang et al. (Wang et al., 2023b) incorporated AU detection to introduce inductive biases that can enhance the generalization of deep MiE recognition.
Self-supervised learning. To enhance the ability to capture details of MiEs, self-supervised learning can reduce noise and irrelevant information in subtle changes of MiEs. Fan et al. (Fan et al., 2023) proposed an MiE analysis method by using self-supervised learning and a symmetric contrastive ViT to capture the symmetrical facial movements in MiEs while mitigating the influence of irrelevant information. Since standard BERT cannot describe MiE features in detail, Nguyen et al. (Nguyen et al., 2023) introduced a self-supervised learning method called micron-BERT that consists of two components: diagonal micro-attention to capture subtle differences between adjacent frames and a patch-of-interest mechanism to emphasize regions within MiEs while mitigating background noise. Das et al. (Das et al., 2021) elaborated a self-supervised approach to learning meaningful dynamic features of MiEs with limited samples as pretext tasks. The pretext model was subsequently adapted for downstream tasks while retaining prior knowledge of facial motion. Motivated by contrastive learning within the momentum contrast framework (He et al., 2020), Wang et al. (Wang and Shang, 2023) utilized a pre-trained 3D-CNN to learn discriminative information embedded in the underlying data structure. To capture subtle and rapid facial movements, the 3D-CNN enhanced temporal dynamics by incorporating a generative adversarial model to remove redundant frames.
Lightweight deep learning. The ever-increasing demand for MiE recognition in daily life drives the development of lightweight and resource-efficient deep MiE recognition methods (Khor et al., 2019; Liong et al., 2019; Ab Razak and Sahran, 2023). Compared with the complex deep models, the lightweight deep model can alleviate the computational demand while maintaining reasonable predictive performance. Khor et al. (Khor et al., 2019) proposed a lightweight dual-stream lightweight deep model comprising two truncated CNNs with distinct input features. By merging different convolutional features from both streams, the proposed method facilitates more effective training and enhances the learning of discriminative features for MiE recognition. Liong et al. (Liong et al., 2019) designed a triple stream 3D-CNN to recognize MiE from three optical flow features, i.e., optical strain, horizontal and vertical optical flow fields. With fewer model parameters than traditional deep models, the proposed method in (Liong et al., 2019) could extract discriminative high-level features to describe the fine-grained MiE details. Shukla et al. (Shukla et al., 2022) introduced a hybrid method combining convolutional layers with LSTM for lightweight MiE recognition. By integrating convolutional and recurrent structures, the proposed model (Shukla et al., 2022) can capture complex spatio-temporal features and address data imbalance issues. To identify global features, Wang et al. (Wang et al., 2023a) proposed a multi-branch attention CNN that consists of region division, multi-branch attention expression learning, and global feature fusion. The proposed multi-branch attention module can extract four distinct features and combine the four extracted features with different weights.
5.2.3. Other promising learning paradigms.
To expand the potential scope of MiE recognition, several promising techniques (Dai and Feng, 2021; Li et al., 2021a; Sharma et al., 2022) have emerged that cater to practical requirements and specific application scenarios. These methods address challenges such as limited datasets and focus on applications across diverse settings.
One approach to train deep models with limited data is to employ meta learning. By training on a variety of tasks, meta-learning enables models to “learn how to learn”, acquiring rapid adaptation capabilities that are valuable in data-scarce scenarios. To address the insufficient and inconsistent labeling in MiE datasets, Dai et al. (Dai and Feng, 2021) introduced a few-shot learning method that transfers knowledge through AUs. The few-shot learning method in (Dai and Feng, 2021) leverages the rich information contained in AUs to enhance the model ability to generalize from limited labeled examples. Wan et al. (Wan et al., 2022) applied model-agnostic meta learning to initialize parameters for 3D CNN training, enabling the model to quickly adapt to new tasks with minimal fine-tuning. Gong et al. (Gong et al., 2023) utilized a meta-learning-based multi-model fusion framework to extract the deep features from frame differences and optical flows. After verification of various feature combinations, the framework demonstrated that the fusion method using MaE and MiE can give distinct information to further recognize MiEs.
Another approach to facilitate insufficient samples in MiE recognition is data generation using a generative adversarial network (GAN) (Goodfellow et al., 2014). The proposal of GAN can be applied in many aspects including image generation (Liao et al., 2022), data augmentation (Tran et al., 2021), super-resolution (Park et al., 2023), and style transfer (Xu et al., 2021a). Recently, MiE studies employed the GAN to generate more available data samples for training robust recognition models. Liong et al. (Liong et al., 2020) used two GAN-based models, namely auxiliary classifier GAN (Odena et al., 2017) and self-attention GAN (Zhang et al., 2019), to generate artificial MiE samples. Li et al. (Li et al., 2021a) proposed an improved GAN for MiE analysis using AU-based multi-label learning to generate sequences that were very similar to the original sequences. This approach ensures high fidelity in the generated data, which is crucial for training robust MiE recognition models. To address the issues of AU annotation caused by a subtle change in the face, Xu et al. (Xu et al., 2021b) introduced a fine-grained AUs modulation to alleviate the noises and handle the symmetry of AUs intensity. Yu et al. (Yu et al., 2021b) provided a GAN-based method, identity-aware and capsule-enhanced generative adversarial network with graph-based reasoning, to conduct controllable MiE synthesis with identity-aware features for recognition. This innovative approach allows for the generation of synthetic MiE sequences that retain subject identity while enhancing recognition performance. Zhou et al. (Zhou et al., 2023) developed a GAN-based approach to enhance the generation of MiE sequences, aiming for more natural and seamless appearances. They first analyzed all AUs presented in prominent MiE datasets. Then, they smoothed the AU matrix extracted from source videos to refine the input data. By leveraging this refined AU information within their GAN framework, they were able to generate MiE sequences that exhibit more realistic and natural transitions. Zhang et al. (Zhang et al., 2023) introduced a facial-prior-guided MiE generation framework for facial motion synthesis. This framework incorporated two key structures: an adaptive weighted prior map and a facial prior module. These components mitigated AU estimation errors and guided motion feature extraction, ensuring smooth and realistic synthesis generation by leveraging facial priors.
The focus of research on MiE recognition has been shifting from controlled laboratory settings to uncontrolled, wild scenarios to better address practical applications. Accurate MiE recognition in such environments presents several challenges, including occlusions, pose variations, varying illuminations, and low-resolution images. To address the challenges and improve the robustness and performance of recognition systems, a variety of techniques have been proposed. One approach to handling low-resolution MiE images in uncontrolled environments is through super-resolution techniques. Sharma et al. (Sharma et al., 2022, 2022) introduced methods based on GAN to perform super-resolution on low-resolution MiE images. This approach transformed low-resolution MiE images to super-resolution, allowing for the extraction of fine-grained textural features useful for downstream tasks. Occlusions represent another significant challenge for MiE recognition in wild conditions. Mao et al. (Mao et al., 2022) alleviated occlusions by generating synthetic datasets featuring various types of occlusions, such as facial masks, glasses, and random region masks, which allow for more thorough analysis under occlusion conditions. Gan et al. (Gan et al., 2022) developed an integrated “spot-and-recognize” framework that incorporated modules for MiE spotting, face alignment, and feature extraction. Building on this foundational work, Gan et al. (Gan et al., 2023) introduced significant enhancements by integrating 3D face reconstruction techniques into the existing framework. Specifically, Gan et al. (Gan et al., 2023) developed methods to generate a 3D face mesh that improves the accuracy of face alignment and the reliability of feature extraction from a 3D perspective. Even under challenging conditions such as varying poses and illuminations, the proposed integration in (Gan et al., 2023) can obtain accurate and consistent MiE recognition.
5.2.4. Potential limitations of different methods on MiE recognition.
Shallow machine learning methods rely heavily on handcrafted descriptors, which require significant domain expertise and are time-consuming and challenging to develop. Feature selection and fusion strategies, while enhancing performance, introduce additional complexity. Ensuring that features are both representative and complementary and without introducing bias or redundancy, is a challenging task.
For MiE recognition, deep learning methods can improve classification performance, but using static images or apex frames as input can result in information loss. Apex frames capture only peak emotional intensity and ignore subtle changes during other time points, affecting the overall understanding of the MiE sequence. Although dynamic images can integrate spatiotemporal information, converting dynamic images into single frames compresses spatiotemporal information. Similarly, using the entire clip as input of the deep MiE recognition model can introduce redundant information.
With MiEs exhibiting minimal changes over short time spans, selecting the appropriate model input is crucial. Transfer learning uses features from related datasets to aid in MiE recognition, but the datasets may not fully align with MiE characteristics. Multi-task learning remains an area for further exploration, particularly in identifying tasks that can strengthen MiE recognition. Lightweight learning methods aim to reduce inference time and storage requirements, but balancing model efficiency with accuracy remains a challenge.
To address the issue of small sample sizes, self-supervised learning leverages pretext tasks to learn meaningful spatiotemporal representations, though designing effective pretext tasks that align with downstream recognition remains a difficulty. Meta-learning and few-shot learning enable models to adapt to new tasks with limited data, but ensuring task diversity for effective training and improving generalization remain significant hurdles. GANs can generate additional training samples, improving model robustness, but the quality of synthetic data and annotation accuracy are critical concerns. Super-resolution techniques, while enhancing image detail, increase computational complexity and present challenges in real-time applications, particularly for large video datasets. Finally, in uncontrolled real-world environments, factors like lighting, occlusions, and pose variations significantly impact MiE recognition, and maintaining consistent performance under these conditions remains a major challenge.
6. Potential Applications of Facial Expression Analysis in IoT
The convergence of facial expression analysis with IoT systems unlocks new chances for intelligent, pervasive and adaptive computing services. By interpreting human emotions, IoT based applications can provide personalized, efficient, and responsive user experiences. In this section, we explore the potential of facial expression analysis in IoT by separately examining the contributions of MaE recognition and holographic MiE analysis.
6.1. MaE and Its Applications in IoT
MaEs play a vital role in interpreting human emotions; and therefore, MaE recognition becomes valuable in various IoT applications. More specifically, the applications of MaE recognition based IoT systems have been extended from purely emotion recognition (Muhammad and Hossain, 2021; Chen et al., 2020a; Yang et al., 2021) to the healthcare industry (Hossain and Muhammad, 2019) and flu tracking (Rahman and Hossain, 2021).
6.1.1. Emotion recognition.
The recent years have seen the rapidly expanding attention on the integration of emotion recognition and intelligent IoT devices. Such integration enhances real-time data analysis, minimal communication overhead, and data protection. Muhammad et al. (Muhammad and Hossain, 2021) designed an edge computing-empowered emotion recognition system that consists of IoT devices, an edge device, and a remote cloud. Pre-processing tasks (such as face detection, cropping, contrast enhancement, and resizing) are performed on IoT devices to alleviate computation and communication expenditure. Then, the edge device leverages the processed images for inference and decision-making and returns the decisions to the IoT devices. Moreover, the edge device downloads a global deep model from the cloud during off-peak hours to minimize latency. The remote cloud manages large-scale training and updates to ensure that the system remains up-to-date. The energy efficiency, lightweight design, and strong performance in experiments of the proposed architecture demonstrate the advantages of integrating emotion recognition with IoT devices.
For the privacy-preserving objective, Chen et al. (Chen et al., 2020a) introduced an IoT-based MaE recognition method. More specifically, due to the imbalanced and limited size of expression datasets, data augmentation is employed to enhance model performance. The IoT-based MaE recognition method ensures that sensitive emotional data is processed locally in order to reduce the risks of privacy leakage. Yang et al. (Yang et al., 2021) proposed a MaE recognition system for human-computer interaction that operates efficiently in real-time. By using a lightweight deep model on IoT devices to recognize emotions through AUs, the proposed system can minimize pre-processing steps and limited data transmission. Consequently, it offers low latency and reduced overhead that satisfies the stringent requirements of real-time applications. Chen et al. (Chen et al., 2022a) developed a system combining face recognition and emotion detection for use in autonomous vehicles. Leveraging IoT devices such as cameras, the system can facial emotions and identities of drivers to enhance takeover effectiveness. Optimized IoT protocols reduce the delay to just 2 milliseconds, while achieving high reliability with an F1 score of 96%.
6.1.2. Healthcare industry.
The application of MaE analysis in the IoT-based healthcare industry has demonstrated a promising potential. By integrating the MaE information with IoT technology, healthcare systems can provide real-time analysis and feedback that are crucial for improving medical services and patient care. For example, Hossain et al. (Hossain and Muhammad, 2019) developed an emotion recognition system for medical facility enhancement. The proposed approach utilized an IoT-based audio-video framework to assist healthcare providers in evaluating services. The system performs audio feature extraction and face detection on edge devices. Then, the remote cloud delivers the well-trained parameters to edge devices for efficient model inference. Such setup allows healthcare providers to gain immediate insights into patient emotions, facilitating more personalized and responsive care. Wang et al. (Wang et al., 2022d) highlights the potential of IoT-based MaE analysis in ensuring worker safety in hazardous industries. More specifically, a deep capsule method based on IoT devices was introduced in (Wang et al., 2022d) to monitor the mental state of miners in underground mining environments. The proposed system collects facial expressions and electroencephalograms to infer mental health states. By integrating the two modalities of data, the system could detect signs of stress and fatigue and thereby reduces the risk of man-made accidents.
6.1.3. Flu tracking.
The application of IoT-based MaE analysis in flu tracking can enhance the management and control of flu spread through real-time monitoring and data analysis. For example, Rahman et al. (Rahman and Hossain, 2021) developed an edge internet of medical things framework to capture relevant signals (e.g., facial expressions and psychological states) during pandemics. By leveraging edge computing, the proposed edge internet of medical things framework can process data locally to generate immediate reports that highlight key insights. Such setup ensures low latency, enhanced security, and robust privacy protection that are critical factors in managing sensitive health information. Moreover, the proposed framework supports a variety of applications, such as sleep analysis, face mask detection, and physiological state analysis during a pandemic.
6.2. MiE and Its Applications in IoT
Since MiE can reveal ground-truth emotions, MiE-based IoT applications are an emerging and promising application scenario. By incorporating MiE analysis, IoT-based frameworks can detect and identify hidden emotions such that the valuable insights for practical applications are provided, e.g., security monitoring (Saeed, 2021; Yildirim et al., 2023; Chung et al., 2014; Shilaskar et al., 2023), interpersonal negotiation (Xiong et al., 2023), and mental disorder detection (Huang, 2021; Das et al., 2021; Gilanie et al., 2022; Li et al., 2024; Chen and Luo, 2023).
6.2.1. Security monitoring.
MiEs  are novel social behavioral biometrics that can be utilized for security monitoring. Due to the natural characteristics, MiEs  provide valuable biometric information that is difficult to pose or fake. Saeed et al. (Saeed, 2021) explored the use of MiEs  as a soft biometric for person recognition. The person recognition utilized three handcrafted features to extract facial and behavioral biometric features from videos containing MiEs  and predicted potential attacks. MiEs  can also be used for deception detection. For instance, Yildirim et al. (Yildirim et al., 2023) made substantial contributions by analyzing facial expressions in video data. They assessed the likelihood of the most prominent MiEs within a given clip or image to achieve accurate deception detection. Widjaja et al. (Widjaja et al., 2023) introduced a deception detection system based on the GRU. This system analyzed meaningful patterns from MiEs in facial videos to identify potential signs of deception. By leveraging the temporal dynamics captured by GRUs, the system can detect subtle changes in facial expressions over time, enhancing its accuracy in identifying deceptive behavior. Shilaskar et al. (Shilaskar et al., 2023) also examined the effectiveness of deception detection using MiEs. Their proposed computer vision-based method evaluates MaEs  and MiEs  captured by a camera to identify specific facial cues associated with deception.
6.2.2. Interpersonal negotiation.
Although emotion can be reflected in facial expressions, it is difficult for someone to capture the emotional changes of another person in time due to the complexity of actual situations. To capture the true emotion and provide timely feedback in negotiation, Xiong et al. (Xiong et al., 2023) introduced a real-time MiE-based emotion recognition system via wearable smart glasses. By collecting and extracting features of the facial video on the front end of smart glasses, users can understand the communicator’s intended information through real-time analysis of MiEs.
6.2.3. Mental disorder detection.
As fleeting information in facial activity, the subtle movements of MiE can reveal psychological states. Huang et al. (Huang, 2021) proposed an elderly depression detection method using the MiEs  based on AU features. Das et al. (Das et al., 2021) presented a method to extract significant MiE muscle movements to predict mental disorders and demonstrated its potential in detecting future stuttering disfluency. Gilanie et al. (Gilanie et al., 2022) proposed a lightweight depression detection method using CNN for real-time applications via digital cameras. Li et al. (Li et al., 2024) achieved promising results in depression recognition by analyzing the MiEs. Chen et al. (Chen and Luo, 2023) developed a method to diagnose concealed depression using MiE combined with ROI and machine learning, offering a low-cost, privacy-preserving solution feasible for self-diagnosis on mobile devices.
7. Challenges and Future Directions
The research on facial expression analysis has sparked numerous notable studies, which resulted in significant advances in both its application and further research. Despite its current prosperity in various domains, there remain several challenges and issues that require attention in the future. Here, we provide three notable research topics and potential research directions in the blow.
7.1. Advancing Facial Expression Datasets
Current datasets used for training effective facial expression analysis systems face significant challenges. In particular, compared to MaE, the available data for MiE is limited. This is due to the subtle and transient nature of MiE, making data collection and annotation more difficult. Most existing MiE datasets are collected in controlled laboratory settings, which may limit their generalizability to real-world applications. Additionally, variations in methodologies and paradigms used for data collection further complicate efforts to standardize and expand these datasets, hindering their scalability and growth.
Annotating MiE data also presents a major challenge. The process requires certified expertise to ensure accuracy and reliability, which significantly increases the time and resource investment needed for dataset creation. While MaE datasets tend to be larger, they also face challenges, primarily due to biases in data collection. Most MaE datasets are derived from publicly available internet sources, leading to inconsistencies in data quality and potential bias.
To address these challenges, future research should focus on several key areas: 1) Unsupervised and Semi-Supervised Learning: These approaches can reduce reliance on large, annotated datasets by enabling models to learn from limited or even unlabeled data; 2) Synthetic Data Generation: Leveraging techniques such as GANs or diffusion models can augment existing datasets with realistic facial expression samples that retain key analytical features; 3) Standardized Protocols for Data Collection: Developing universal standards for collecting and annotating facial expression data across diverse settings can improve dataset consistency and interoperability. This includes protocols designed for “wild” data collection, such as using wearable devices or ubiquitous cameras. Data captured by IoT devices can provide a more diverse and naturalistic representation of facial expressions.
7.2. Adaptation to Diverse Cultural Contexts
Facial expression analysis systems often rely on the assumption of universal emotional expressions. However, cultural differences play a significant role in how emotions are expressed and interpreted. While MiEs are considered universal indicators of hidden or suppressed emotions, their occurrence may vary across cultures. On the other hand, MaEs are strongly influenced by culturally learned norms that dictate the appropriateness of expressing certain emotions in specific contexts. To effectively adapt facial expression analysis systems to diverse cultural settings, datasets must include both MiEs and MaEs from a variety of cultural contexts. To improve accuracy and relevance, models should incorporate contextual information, such as the interaction environment and cultural norms.
This cultural adaptability becomes especially important in IoT systems, as devices like smart home assistants, surveillance cameras, and wearable technologies are deployed globally. For instance, the facial expression details that smart assistants in different cultural settings need to focus on may vary significantly. To achieve cultural adaptability in IoT systems, it is imperative to collect culturally diverse datasets, including expressions from different geographic regions, age groups, and social contexts.
MiE datasets should capture subtle, transient expressions across various cultures, while MaE datasets must reflect culturally influenced emotional intensity and duration. Furthermore, pre-trained models can be fine-tuned using culturally specific datasets through transfer learning. This approach enables IoT devices to adapt to regional needs without requiring complete retraining from scratch. Additionally, IoT devices should integrate contextual understanding into their emotion recognition pipelines. For example, wearable devices like smart glasses can combine facial expression data with environmental cues such as social setting or cultural event, to provide culturally sensitive feedback or assistance.
7.3. Data Privacy and Protection
Data privacy and security are critical concerns in the integration of facial expression analysis into IoT systems. Facial expression data often consists of images or videos of individuals, which can include sensitive biometric information. The collection, transmission, and storage of such data from distributed IoT devices can introduce significant privacy risks. While some datasets are collected under strict informed consent protocols, many IoT-derived datasets remain unregulated. The reliance on wireless communication protocols in IoT devices increases vulnerability to interception and hacking. Weak encryption and unsecured channels can expose sensitive facial data during transmission. Additionally, storing facial data in centralized servers or cloud-based systems raises the risk of data breaches. A lack of a unified global privacy protection framework further exacerbates these concerns. While regulations like the European Union General Data Protection Regulation set standards for biometric data protection, the enforcement of these regulations varies by country, complicating global compliance.
Synthetic facial expression data has emerged as a promising solution to address these privacy concerns. Using GANs, synthetic data can be generated to supplement real datasets or to enhance model performance. Moreover, with advancements in generative AI, large-scale text-to-video models (e.g., Sora) now offer the ability to create realistic facial expression data, providing a potential avenue for safeguarding privacy in future applications.
Federated learning is another promising technique for enhancing privacy in facial expression analysis. Unlike traditional centralized machine learning approaches, federated learning enables multiple devices to collaboratively train models while keeping sensitive data localized and without sharing raw data. This decentralized approach ensures that facial data never leaves the user’s device, reducing the risk of unauthorized access and data leakage. By only sharing aggregated model updates with a central server, federated learning prevents the reconstruction of individual user’s data. However, this approach requires efficient algorithms that can run on resource-constrained IoT devices, so future research should focus on optimizing federated learning techniques to balance performance, privacy, and energy efficiency.
7.4. Real-Time Processing
Real-time processing is a critical requirement for facial expression analysis in IoT applications. For instance, emotion-driven smart assistants and healthcare monitoring systems depend on real-time performance to function effectively. However, achieving efficient, reliable, and scalable real-time facial expression analysis remains a significant challenge. These challenges arise from the computational complexity of facial expression analysis algorithms, the resource constraints of IoT devices, and the difficulties associated with deploying these systems in the wild.
One of the primary challenges is the high computational cost of facial expression analysis algorithms. Tasks such as feature extraction, expression classification, and temporal analysis demand substantial processing power. Deep learning models, which are commonly used for these tasks, involve millions of parameters and require significant computational resources. IoT devices, however, are often resource-constrained in terms of processing power, memory, and battery life, making it difficult to meet the real-time processing demands of these algorithms.
Latency is another critical issue. IoT systems that rely on centralized cloud servers for processing can suffer from delays due to data transmission and network bottlenecks. These delays can undermine the responsiveness of applications, particularly in time-sensitive areas such as healthcare. Ensuring low-latency communication and processing is essential for effective real-time performance.
Edge computing presents a promising solution to these challenges. By processing data locally on IoT devices rather than sending it to centralized servers, edge computing reduces latency and conserves bandwidth. However, implementing facial expression analysis at the edge requires efficient hardware accelerators such as GPUs or specialized AI chips. Additionally, techniques like model quantization, pruning, and knowledge distillation can help optimize facial analysis models to reduce computational requirements while maintaining accuracy. Efficient data transmission protocols are also crucial for real-time performance, particularly for IoT networks with limited bandwidth and strict latency requirements.
7.5. Comprehensive Analysis
Comprehensive facial expression analysis involves multi-modal approaches that extend beyond the traditional focus on facial muscle movements. Recent studies have introduced various modalities such as electromyography (Kim et al., 2022), word (Wahid et al., 2023), electroencephalogram (Wang et al., 2023c; Zhao et al., 2024) and other psychological signals (Saffaryazdi et al., 2022) can also reveal the emotions. Developing a multi-modal analysis of facial expressions could take advantage of complementary information and obtain robust features of emotions. Furthermore, several datasets combined with physiological signals have been proposed. For example, Chen et al. (Chen et al., 2019) introduced the micro-gesture dataset for hidden emotion recognition. Zou et al. (Zou et al., 2022) proposed a facial expression dataset with synchronized physiological signals. More valuable research can explore the intrinsic relationship between facial expressions and other signals in the future. In addition to applying multi-modal expression analysis methods to IoT, IoT needs to consider multi-tasks for collaborative analysis.
In the context of the IoT, multi-modal expression analysis should also address the challenges of multi-tasking. IoT devices often need to process data from diverse sensors and perform various analyses simultaneously. This necessitates the development of efficient multi-task learning models to enhance system performance and responsiveness. For example, a smart home system might simultaneously analyze data from security cameras, environmental sensors, and facial expressions to enable collaborative decision-making.
Future directions for facial expression analysis in IoT should prioritize the integration of multi-modal and multi-task capabilities. By incorporating various physiological signals, researchers can develop more accurate and comprehensive emotion recognition systems. In addition, by improving the efficiency of multi-task learning models, IoT devices can maintain high performance and responsiveness while executing complex tasks. This will enable IoT systems to better understand and respond to emotional demands of users to provide more intelligent and human-centric services.
8. Conclusions
Facial expression analysis is receiving perpetually rising attention to understand the rich emotional information. As a common type of facial expression, the MaE analysis is now used in various wild scenarios. In contrast, the MiE analysis with the characteristics of unconsciousness and subtlety has the potential to reveal the genuine emotions of each individual. In this article, we have provided the taxonomy of current facial expression analysis that includes MaE and MiE analysis. We have comprehensively reviewed the state-of-the-art facial expression analysis methods and discussed their corresponding limitations. By reviewing the current applications of MaE and MiE in IoT systems, we have discussed insights and potential directions for the development and implementation of facial expression analysis in future IoT systems. We expect that our work can serve as a valuable resource for researchers and practitioners in facial expression analysis by providing fundamental research resources and a comprehensive research paradigm.